### v1

# The polytope of all $q$-rank functions

Abstract.
A $q$-rank function is a real-valued function defined on the subspace lattice that is non-negative, upper bounded by the dimension function, non-drecreasing, and satisfies the submodularity law. Each such function corresponds to the rank function of a $q$-polymatroid. In this paper, we identify these functions with points in a polytope. We show that this polytope contains no interior lattice points, implying that the points corresponding to $q$-matroids are among its vertices. We investigate several properties of convex combinations of two lattice points in this polytope, particularly in terms of independence, flats, and cyclic flats. Special attention is given to the convex combinations of paving and uniform $q$-matroids.
Keywords. polytopes, $q$-rank function, $q$-polymatroids, rank-metric codes.
1. Introduction
The concept of matroid abstracts the notion of independence in vector spaces and graphs. Matroids have been extended in various directions, including polymatroids [7], valuated matroids [6], and, more recently, latroids [25, 17] and $q$-analogues such as $q$-matroids [20], $q$-polymatroids [16, 24], and $\mathcal{L}$-polymatroid [1, 4].
In the classical setting, let $S$ be a finite set of cardinality $n$. The rank function of a polymatroid is a real-valued function defined on all the subsets of $S$, which is non-decreasing, submodular, and upper bounded by the cardinality function. The cone of all such polymatroid rank functions has been studied in [22], particularly to characterize its extremal rays and extremal polymatroids. Further developments appear in [10, 21], and analogous investigations have been conducted for supermodular functions; see [18].
In this paper, we focus on $q$-rank functions, i.e. rank functions of $q$-polymatroids. Let $\mathbb{F}_{q}$ be the finite field of order $q$, let $E$ be an $n$-dimensional vector space over $\mathbb{F}_{q}$ and let $\mathcal{L}(E)$ be the lattice of subspaces of $E$ ordered with inclusion. A $q$-rank function is defined on $\mathcal{L}(E)$, and it is real-valued, non-negative, upper bounded by the dimension function, non-decreasing and submodular. If, in particular, such a function is integer-valued, then it is the rank function of a $q$-matroid. We define the polytope of all $q$-rank functions $\mathcal{P}_{q}^{n}\subseteq\mathbb{R}^{|\mathcal{L}(E)|}$ as the polytope defined by the linear inequalities given by the above properties. More specifically, up to ordering the spaces in $\mathcal{L}(E)$, we have that a point in $(p_{X})_{X\in\mathcal{L}(E)}\in\mathcal{P}_{q}^{n}$ has coordinates indexed by the spaces in $\mathcal{L}(E)$ and the entries are precisely the values that a $q$-rank function takes in each $X\in\mathcal{L}(E)$. Hence, the integer points in this polytope correspond precisely to $q$-matroids. We show that $\mathcal{P}_{q}^{n}$ has dimension $|\mathcal{L}(E)|-1$; see Proposition 3.5. However, as this dimension grows rapidly with $n$, analyzing its geometry becomes challenging. One of our main results in this direction is that $\mathcal{P}_{q}^{n}$ contains no integer points (lattice points) in its interior; see Theorem 3.8. In particular, this means that all $q$-matroids correspond to the vertices with integer coordinates. It is clear that as soon as $n$ grows, the number of vertices also increases. Being this a polytope described by constraints with rational coefficients, it is a rational polytope, i.e. all its vertices have rational coordinates.
We introduce convex combinations of points in $\mathcal{P}_{q}^{n}$ as an operation on $q$-polymatroids. Beyond establishing general results on the flats of such combinations (see Section 4.1), we focus in particular on the convex combination with rational coefficients of $q$-matroids. We investigate how key structural elements, such as cyclic spaces and $\mu$-independent spaces (where $\mu$ is a denominator of the resulting $q$-polymatroid), behave under this operation. Specifically, we express these features of the resulting $q$-polymatroid in terms of the cyclic spaces and independent spaces of the original $q$-matroids. We specialize our analysis to convex combinations involving two important families of $q$-matroids: paving $q$-matroids and uniform $q$-matroids. For paving $q$-matroids, we focus particularly on a construction introduced in [12]. In this context, we introduce the notion of $\mu$-pavingness for $q$-polymatroids and show that the convex combination of two paving $q$-matroids, built using this special construction, yields a $\mu$-paving $q$-polymatroid, where the denominator $\mu$ is not necessarily principal; see Theorem 5.6. We further characterize important structural features of the resulting $q$-polymatroid: its flats (Theorem 5.11), cyclic spaces (Theorem 5.12), and cyclic flats (Corollary 5.13). For the convex combination of two uniform $q$-matroids, we provide characterizations of $\mu$-independent spaces, under specific constraints (see Theorems 6.1 and 6.2), and cyclic flats (see Proposition 6.8). These ideas are then extended to the convex combination of $n-2$ uniform $q$-matroids, yielding further results on $\mu$-independence (see Theorem 6.5).
Another well-studied concept in matroid theory is the characteristic polynomial. Generalizations of this invariant have been proposed for $q$-matroids, integer $q$-polymatroids, and $\mathcal{L}$-matroids; see [19, 3, 1]. In this work, we introduce a broader notion of the characteristic polynomial, tailored for $q$-polymatroids whose $q$-rank functions take rational values. Specifically, we define an invariant, the characteristic Puiseux polynomial, associated with such $q$-polymatroids. This invariant is not a polynomial in the classical sense: it is a truncated Puiseux series, meaning a finite linear combination of monomials whose exponents may be rational (and possibly negative). Despite this departure from classical polynomials, we adopt the name characteristic Puiseux polynomial to emphasize its structural role and its connection to existing generalizations in matroid theory. We show that the characteristic Puiseux polynomial of the convex combination with rational coefficients of two paving $q$-matroids obtained from the construction in [12] not only can be expressed in terms of their characteristic polynomials, but only one of them is enough; see Theorem 7.4.
One of the drawbacks of the polytope of all $q$-rank function is that we are not able to distinguish representable $q$-polymatroids. This difficulty is due to the lack of an axiomatic definition that applies only for the representable case. However, we illustrate some examples of convex combination of special representable $q$-polymatroids arising from rank-metric codes; see Theorem 8.8.
The rest of the paper is organized as follows. In Section 2, we provide the basic notions on $q$-polymatroids. In Section 3, we introduce the polytope of all $q$-rank functions and initiate the study of its geometric properties. In Section 4, we define the convex combination of $q$-polymatroids and provide its first properties. In Section 5, we characterize the convex combination with rational coefficients of some paving $q$-matroids. In Section 6, we consider the convex combination of uniform $q$-matroids. In Section 7, we introduce the Puiseux characteristic polynomial for rational $q$-rank functions and we study the one of the convex combination of paving $q$-matroids. In Section 8, we provide some examples of convex combination of representable $q$-polymatroids. We conclude with some open questions and future research directions in Section 9.
Acknowledgements
The authors are thankful to Lukas Kühne and Jan Stricker for fruitful discussions. Moreover, they are grateful to Bernd Sturmfels for inspiring this work. The second author would like to especially thank Lukas Kühne for his supervision and guidance throughout the research process. The second author is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – SFB-TRR 358/1 2023 – 491392403 and SPP 2458 – 539866293.
2. Preliminaries
In this section, we outline some basic facts and definitions of $q$-polymatroids. Throughout this paper we will make use of the following notation.
Notation 2.1.
Let $q$ be a prime power and let $\mathbb{F}_{q}$ be the finite field with $q$ elements. Let $n\geq 2$ be a positive integer and $E=\mathbb{F}_{q}^{n}$. For a space $X$, we denote by $\mathcal{L}(X)$ the collection of subspaces of $X$. For $0\leq i\leq\dim(X)$, we denote by $\mathcal{L}(X)_{\leq i}$ and $\mathcal{L}(X)_{\geq i}$ the collections of all subspaces of $X$ of dimension at most $i$ and at least $i$, respectively, while we denote by $\mathcal{L}(X)_{i}$ the collection of all $i$-dimensional subspaces of $X$. Further, we denote by $\mathcal{L}(X)^{\ast}$ the collection $\mathcal{L}(X)\setminus\langle 0\rangle$. For ease of notation, given a space $X$, we denote by $\mathcal{H}(X)$ the set of all codimension-one subspaces of $X$. We write $X\leq Y$ to indicate that $X$ is a subspace of $Y$. If a subspace is to be understood as being one-dimensional, we represent it by a lowercase letter, so for instance, $x\leq X$ means that $x$ is a one-dimensional subspace of $X$. Finally, we set $T:=|\mathcal{L}(E)|$.
The first results on $q$-polymatroids and their relation to coding theory can be read in [16] and [24]. The following definition is from [16, Def.4.1].
Definition 2.2.
A function $\rho:\mathcal{L}(E)\longrightarrow\mathbb{R}$ is a $q$-rank function if it satisfies the following axioms.
  • (R1)
    Boundedness: $0\leq\rho(A)\leq\dim(A)$, for all $A\in\mathcal{L}(E)$.
  • (R2)
    Monotonicity: $A\leq B\Rightarrow\rho(A)\leq\rho(B)$, for all $A,B\in\mathcal{L}(E)$.
  • (R3)
    Submodularity: $\rho(A\cap B)+\rho(A+B)\leq\rho(A)+\rho(B)$, for all $A,B\in\mathcal{L}(E)$.
A $q$-polymatroid is a pair $\mathcal{M}=(\mathcal{L}(E),\rho)$ for which $\rho$ is a $q$-rank function.
The value $\rho(E)$ is called the rank of $\mathcal{M}$.
The definition of $q$-polymatroid appeared in different forms in the literature. In [11, Def. 2.1] the same definition occurs with the only difference that the rank function takes only rational values. In this case we talk about rational $q$-polymatroids. Fundamental concepts regarding rational $q$-polymatroids have been investigated in [12] and [11]. A number $\mu\in\mathbb{Q}_{\geq 0}$ is called a denominator of $\rho$ and $\mathcal{M}$ if $\mu\rho(X)\in\mathbb{N}_{0}$ for all $X\in\mathcal{L}(E)$. The smallest denominator is called the principal denominator. A $q$-polymatroid with principal denominator $1$ (i.e., $\rho(X)\in\mathbb{N}_{0}$ for all $X$) is a $q$-matroid.
Next, for any $r\in\mathbb{N}$ a $(q,r)$-polymatroid as in [24, Def. 2] by Shiromoto, [9, Def. 1] and [3, Def. 1] can be turned into a rational $q$-polymatroid with denominator $r$ by dividing the rank function by $r$. Conversely, given a rational $q$-polymatroid with denominator $\mu$, then $(\mathcal{L}(E),\mu\rho)$ is a $(q,\lceil\mu\rceil)$-polymatroid. Finally, in [1, Def. 2.12] and [4, Def. 2.2] an $\mathcal{L}$-polymatroid is defined as a direct extension of the definition of the previous $q$-polymatroid.
We will use the following definition of independent spaces derived in [11, Def. 3.1].
Definition 2.3.
Let $\mathcal{M}=(\mathcal{L}(E),\rho)$ be a rational $q$-polymatroid with denominator $\mu$, not necessarily principal. A space $I\in\mathcal{L}(E)$ is called $\mu$-independent if
$\rho(J)\geq\frac{\dim(J)}{\mu},$
[…]
In [11, Def. 3.8], the authors define the $\mu$-bases of a space $V\in\mathcal{L}(E)$ as the inclusion-maximal $\mu$-independent subspaces of $V$. Furthermore, they show that all the $\mu$-bases of $V$ have the same rank value, which is equal to the rank of $V$.
In case of $\mathcal{M}$ being a $q$-matroid, the notion of strong independence coincides with the classical one. In particular, we say that a subspace $A\in\mathcal{L}(E)$ is called an independent space of the $q$-matroid $\mathcal{M}$ if $\rho(A)=\dim A$. If $x\leq E$ and $r(x)=0$, then $x$ is called loop of $\mathcal{M}$. By subodularity, it follows that if $x,y\leq E$ are loops, also $x+y$ has to be a loop. Hence, it follows that there is a space containing all the loops. We call such a space loop space. A subspace of $E$ that is not an independent space of $\mathcal{M}$ is called a dependent space of $\mathcal{M}$. We call $C\in\mathcal{L}(E)$ a circuit if it is itself a dependent space and every proper subspace of $C$ is independent.
We finally define flats and cyclic spaces of a $q$-polymatroid $\mathcal{M}=(\mathcal{L}(E),\rho)$.
Definition 2.4.
Let $X\in\mathcal{L}(E)$. We say that $X$ is a flat of $\mathcal{M}$ if $\rho(X)<\rho(X+v)$ for all one dimensional spaces $v\in\mathcal{L}(E)_{1}\setminus\mathcal{L}(X)_{1}$.
The following definition comes from [4, Def. 4.2].
Definition 2.5.
Let $X\in\mathcal{L}(E)$. We say that $X$ is cyclic if for all $H\in\mathcal{H}(X)$ one of the following is satisfied:
  1. (1)
    $\rho(X)=\rho(H)$ or
  2. (2)
    $0<\rho(X)-\rho(H)$ and there exists $a\in\mathcal{L}(X)_{1}\setminus\mathcal{L}(H)_{1}$ such that $\rho(X)-\rho(H)<\rho(a)$.
Furthermore, we call a space $X\in\mathcal{L}(E)$ a cyclic flat of a $q$-polymatroid $\mathcal{M}$ if its both a flat and a cyclic space of $\mathcal{M}$. We denote by $\mathcal{F}(\mathcal{M}),\;\mathcal{O}(\mathcal{M})$ and $\mathcal{Z}(\mathcal{M})$ the collections of flats, cyclic spaces and cyclic flats of a $q$-polymatroid $\mathcal{M}$.
Note that for $q$-matroids the notion of flats is identical, while, for cyclic spaces, the condition (2) is superfluous.
Next we recall two well-known families of $q$-matroids, namely uniform $q$-matroids ([20]) and paving $q$-matroids ([13]), which we will use in the next sections.
Definition 2.6.
Let $0\leq k\leq n$. For each $V\in\mathcal{L}(E)$, define $\rho(V):=\min\{k,\dim(V)\}$. Then $(\mathbb{F}_{q}^{n},\rho)$ is a $q$-matroid. It is called the uniform $q$-matroid on $\mathbb{F}_{q}^{n}$ of rank $k$ and is denoted by $\mathcal{U}_{k,n}(q)$.
Definition 2.7.
A $q$-matroid is said to be paving if for every circuit $C$ we have $\dim(C)\geq\rho(E)$.
3. The polytope of all q-rank functions
In this section, we introduce the polytope of all $q$-rank functions and we describe some of its properties. In order to do this, we endow $\mathcal{L}(E)$ with a linear order in the following way
$\langle 0\rangle\preceq\mathcal{L}(E)_{1}\preceq\mathcal{L}(E)_{2}\preceq\ldots\preceq\mathcal{L}(E)_{n-1}\preceq E,$
where $\mathcal{L}(E)_{i}$ is also endowed with a linear order for each $11\leq i\leq n-$ 1. For every $0\leq\ell\leq n$, the number of $i$-dimensional subspaces of $E$ is denoted by the Gaussian binomial coefficient
$\left[\begin{matrix}n\\ \ell\end{matrix}\right]_{q}=\frac{(q^{n}-1)(q^{n-1}-1)\cdots(q^{n-\ell+1}-1)}{(q-1)(q^{2}-1)\cdots(q^{\ell}-1)}.$
Hence, the size of $\mathcal{L}(E)$ is $T=\sum\limits_{i=0}^{n}\left[\begin{matrix}n\\ i\end{matrix}\right]_{q}$.
Now we consider the vector space $\mathbb{R}^{T}$, which contains vectors of the form $v=(v_{X})_{X\in\mathcal{L}(E)}$.
The three rank axioms from Definition 2.2 define a polyhedron in $\mathbb{R}^{T}$, in which each point $v=(v_{X})_{X\in\mathcal{L}(E)}$ represents a $q$-rank function $\rho$ on $\mathcal{L}(E)$. In other words, each such vector $v$ is the image vector of some $q$-rank function $\rho$, i.e., $v=(\rho(X))_{X\in\mathcal{L}(E)}$. Moreover, by the first rank axiom we have that every coordinate direction is non-negative and bounded from above, therefore this polyhedron is in fact a polytope in $\mathbb{R}^{T}$ that lies completely in $\mathbb{R}_{\geq 0}^{T}$.
Definition 3.1.
We define the polytope of all $q$-rank functions on $E$ or $q$-rank polytope on $E$ as
[…]
$=$ $\{v\in\mathbb{R}_{\geq 0}^{T}\;|\;v\text{ satisfies the constraints }(*)\},$
where the constraints ()(*)( ∗ ) are the following:
  1. (1)
    $v_{X}\leq\dim(X)$, for all $X\in\mathcal{L}(E)$.
  2. (2)
    $v_{X}-v_{Y}\leq 0$, for all $X,Y\in\mathcal{L}(E)$ such that $X\leq Y$.
  3. (3)
    $v_{X+Y}+v_{X\cap Y}-v_{X}-v_{Y}\leq 0$, for all $X,Y\in\mathcal{L}(E)$.
We refer to these inequalities as of type 1, type 2 or type 3.
Remark 3.2.
The $q$-rank polytope $\mathcal{P}_{q}^{n}$ contains all the possible $q$-rank functions. The points of $\mathcal{P}_{q}^{n}$ with integer coordinates (lattice points) correspond to $q$-matroid rank functions.
Example 3.3.
Consider $E=\mathbb{F}_{2}^{2}$ and $\mathcal{L}(E)=\{\langle 0\rangle\preceq\langle 01\rangle\preceq\langle 10\rangle\preceq\langle 11\rangle\preceq E\}$. Then the constraints describing $\mathcal{P}_{2}^{2}$ are given as follows:
Type $1$ Type $2$ Type $3$
$v_{1}=0$ $v_{2}-v_{5}\leq 0$ $v_{5}-v_{2}-v_{3}\leq 0$
$v_{2}\leq 1$ $v_{3}-v_{5}\leq 0$ $v_{5}-v_{2}-v_{4}\leq 0$
$v_{3}\leq 1$ $v_{4}-v_{5}\leq 0$ $v_{5}-v_{3}-v_{4}\leq 0$
$v_{4}\leq 1$
$v_{5}\leq 2$
Note that we omitted all inequalities of type $2$ and $3$ involving subspaces pairs $(\langle 0\rangle,X)$ and also all inequalities of type $3$ involving subspaces pairs $(X,E)$, since they are redundant. Using these inequalities we can determine the $\mathcal{H}$-representation of $\mathcal{P}_{2}^{2}\subseteq\mathbb{R}^{5}$, given by $Av\leq b$, for
$A=\begin{pmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\\ 0&1&0&0&-1\\ 0&0&1&0&-1\\ 0&0&0&1&-1\\ 0&-1&-1&0&1\\ 0&-1&0&-1&1\\ 0&0&-1&-1&1\end{pmatrix}\in\mathbb{R}^{10\times 5}\quad\text{and}\quad b=\begin{pmatrix}0\\ 1\\ 1\\ 1\\ 2\\ 0\\ 0\\ 0\\ 0\\ 0\\ 0\end{pmatrix}\in\mathbb{R}^{10}.$
By making use of the computer algebra system OSCAR [23], we can compute all the lattice points of $\mathcal{P}_{2}^{2}$ and obtain the following six points:
$(0,0,0,0,0),(0,1,1,1,1),(0,1,1,1,2),(0,1,1,0,1),(0,1,0,1,1),(0,0,1,1,1).$ .
Each of these points corresponds to one of the six $q$-matroids on $\mathbb{F}_{2}^{2}$, namely the first three correspond to $\mathcal{U}_{0,2}(2),\mathcal{U}_{1,2}(2),\mathcal{U}_{2,2}(2)$ respectively and the last three to the $q$-matroids of rank $1$ having two bases elements. Moreover we get that these points form the vertex set of $\mathcal{P}_{2}^{2}$.
Remark 3.4.
As Example 3.3 illustrated, some inequalities describing $\mathcal{P}_{q}^{n}$ are redundant. In particular, we can make the following observations.
  1. (a)
    For inequalities of type $2$, we only have to consider pairs of subspaces $X,Y\in\mathcal{L}(E)^{\ast}$ such that $X\lessdot Y$ in the lattice $\mathcal{L}(E)$, i.e. $Y$ covers $X$. This is because for $X\lessdot Z\lessdot Y$ we have inequalities $v_{X}\leq v_{Z}$ and $v_{Z}\leq v_{Y}$ that already imply $v_{X}\leq v_{Y}$.
  2. (b)
    For inequalities of type $3$, we do not have to consider pairs $X,Y\in\mathcal{L}(E)$ such that $X\leq Y$ (or $Y\leq X$). This is because we would get $v_{X+Y}+v_{X\cap Y}=v_{X}+v_{Y}$, since $X+Y=Y$ and $X\cap Y=X$ (or $X+Y=X$ and $X\cap Y=Y$).
Another important observation is that we always have the equality $v_{\langle 0\rangle}=0$ by the first rank axiom. Thus, the $q$-rank polytope is always contained in $\mathbb{R}^{T-1}$. We call the polytope obtained from $\mathcal{P}_{q}^{n}$ by removing the first coordinate, the reduced $q$-rank polytope and denote it by $\bar{\mathcal{P}}_{q}^{n}$. Note that the reduced rank polytope has the same algebraic and geometric properties as the $q$-rank polytope.
Throughout this paper we will always indicate if we currently work in $\mathcal{P}_{q}^{n}$ or $\bar{\mathcal{P}}_{q}^{n}$ respectively.
Now we will determine some geometric properties of $\bar{\mathcal{P}}_{q}^{n}$, especially by computing its dimension and lattice points. First of all, we have the following straightforward result.
Proposition 3.5.
$\bar{\mathcal{P}}_{q}^{n}$ has dimension $T-1$.
Proof.
Let $\rho$ be a real-valued function on $\mathcal{L}(E)$ defined by $\rho(X)=\frac{\dim(X)}{\dim(X)+1}$ for every $X\in\mathcal{L}(E)$. Then $\rho$ is a $q$-rank function. In fact, (R1) and (R2) are clearly satisfied with strict inequalities. In order to show (R3), consider two spaces $X,Y\in\mathcal{L}(E)$, such that $X\nleq Y$ or $Y\nleq X$. For ease of notation, let $x=\dim(X)$, $y=\dim(Y)$, $z=\dim(X\cap Y)$ and $t=\dim(X+Y)=x+y-z$. In particular, $0<z<\min\{x,y\}$ and $x,y<t$. We show that $\frac{t}{t+1}+\frac{z}{z+1}<\frac{x}{x+1}+\frac{y}{y+1}.$ We then have the following
$\begin{array}[]{cccc}&\frac{t}{t+1}+\frac{z}{z+1}&<&\frac{x}{x+1}+\frac{y}{y+1}\\ \Leftrightarrow&\frac{1}{t+1}+\frac{1}{z+1}&>&\frac{1}{x+1}+\frac{1}{y+1}\\ \Leftrightarrow&\frac{x+y+2}{(x+y-z+1)(z+1)}&>&\frac{x+y+2}{(x+1)(y+1)}\\ \Leftrightarrow&(x+y-z+1)(z+1)&<&(x+1)(y+1)\\ \Leftrightarrow&(x-z)(z-y)&<&0.\\ \end{array}$
The last inequality is always satisfied, so also (R3) is satisfied with strict inequality. Then, the correspondent point $P_{\rho}$ in $\bar{\mathcal{P}}_{q}^{n}$ is in the interior of $\bar{\mathcal{P}}_{q}^{n}$. Hence, we conclude that the dimension of $\bar{\mathcal{P}}_{q}^{n}$ is $T-1$. ∎
Determining in general the faces of $\bar{\mathcal{P}}_{q}^{n}$ is not an easy task. In the following example, we examine the case of $\bar{\mathcal{P}}_{2}^{2}$, which is the smallest one we can deal with.
Example 3.6.
Let $E=\mathbb{F}_{2}^{2}$. First of all, we have that $\bar{\mathcal{P}}_{2}^{2}\subset\mathbb{R}^{4}$ and, by using the lexicographic linear order, we have $\langle 01\rangle\preceq\langle 10\rangle\preceq\langle 11\rangle\prec E$. As already established, the dimension of $\bar{\mathcal{P}}_{2}^{2}$ is $4$. A face of a polytope is any intersection of the polytope with a supporting hyperplane, which is a hyperplane non trivially intersecting the polytope whose closed half-spaces contain the polytope. By using the computer algebra system OSCAR [23] we compute that there exist $6$ many 00-dimensional faces, $15$ many $1$-dimensional faces, $19$ many $2$-dimensional faces and $9$ many $3$-dimensional faces of $\bar{\mathcal{P}}_{2}^{2}$. We illustrate some of them.
00-faces. For all $v\in\bar{\mathcal{P}}_{2}^{2}$ we have that $v_{1}+v_{2}+v_{3}+v_{4}\leq 5$. Then
$F_{0}=\bar{\mathcal{P}}_{2}^{2}\cap\{v\in\mathbb{R}_{\geq 0}^{4}\;|\;v_{1}+v_{2}+v_{3}+v_{4}=5\}=\{(1,1,1,2)\}$
is 00-dimensional and yields us a vertex of $\bar{\mathcal{P}}_{2}^{2}$. Moreover, if we consider the intersections of $\bar{\mathcal{P}}_{2}^{2}$ with the hyperplanes $\{v\in\mathbb{R}_{\geq 0}^{4}\;|\;v_{2}+v_{3}+v_{4}=4\}$ and $\{v\in\mathbb{R}_{\geq 0}^{4}\;|\;v_{1}+v_{3}+v_{4}=4\}$ we get different descriptions of $F_{0}$. We also have that $v_{i}+v_{4}\leq 3$, for $i=1,2,3$. The intersections of $\bar{\mathcal{P}}_{2}^{2}$ with the hyperplanes $H_{i}=\{v\in\mathbb{R}_{\geq 0}^{4}\;|\;v_{i}+v_{4}=3\}$, gives $F_{0}$, for each $i=1,2,3$. Finally, we have that $v_{4}\leq 2$. If we consider the intersection of $\bar{\mathcal{P}}_{2}^{2}$ with the hyperplane $\{v\in\mathbb{R}_{\geq 0}^{4}\;|\;v_{4}=2\}$, we get again the same face $F_{0}$.
$1$-faces. For all $v\in\bar{\mathcal{P}}_{2}^{2}$ we have that $v_{1}+v_{2}+v_{3}\leq 3$. Then
$F_{1}$ $=\bar{\mathcal{P}}_{2}^{2}\cap\{v\in\mathbb{R}_{\geq 0}^{4}\;|\;v_{1}+v_{2}+v_{3}=3\}$
[…]
is $1$-dimensional and contains only two $q$-matroids, namely $\mathcal{U}_{2,2}(2)$ and $\mathcal{U}_{1,2}(2)$.
$2$-faces. For all $v\in\bar{\mathcal{P}}_{2}^{2}$ we have that $v_{1}+v_{2}\leq 2$. Then
$F_{2,1}$ $=\bar{\mathcal{P}}_{2}^{2}\cap\{v\in\mathbb{R}_{\geq 0}^{4}\;|\;v_{1}+v_{2}=2\}$
[…]
is $2$-dimensional and contains the lattice points $(1,1,1,2),(1,1,1,1),(1,1,0,1)$, which correspond to $\mathcal{U}_{2,2}(2)$, $\mathcal{U}_{1,2}(2)$ and the $q$-matroid, having loopspace $\langle 11\rangle$, respectively. Similarly, we can consider
$F_{2,2}$ $=\bar{\mathcal{P}}_{2}^{2}\cap\{v\in\mathbb{R}_{\geq 0}^{4}\;|\;v_{1}+v_{3}=2\}$
[…]
and
$F_{2,3}$ $=\bar{\mathcal{P}}_{2}^{2}\cap\{v\in\mathbb{R}_{\geq 0}^{4}\;|\;v_{2}+v_{3}=2\}$
[…]
We get that $F_{2,2}$ and $F_{2,3}$ are also $2$-faces containing both $\mathcal{U}_{1,2}(2)$ and $\mathcal{U}_{2,2}(2)$. $F_{2,2}$ also contains the $q$-matroid with loopspace $\langle 01\rangle$ and $F_{2,3}$ contains the $q$-matroid with loopspace $\langle 10\rangle$ respectively.
$3$-faces. For all $v\in\bar{\mathcal{P}}_{2}^{2}$ we have that $v_{i}\leq 1$, for $i=1,2,3$. Then for each $i=1,2,3$ we have that
$F_{3,i}=\bar{\mathcal{P}}_{2}^{2}\cap\{v\in\mathbb{R}_{\geq 0}^{4}\;|\;v_{i}=1\}$
is $3$-dimensional. $F_{3,1}$ contains the lattice points $(1,1,1,2),(1,1,1,1),(1,1,0,1),(1,0,1,1)$. $F_{3,2}$ contains the lattice points $(1,1,1,2),(0,1,1,1),(1,1,0,1),(1,1,1,1)$ and $F_{3,3}$ contains the lattice points $(1,1,1,2),(0,1,1,1),(1,0,1,1),(1,1,1,1)$.
Finally recall that $\bar{\mathcal{P}}_{2}^{2}$ has only six lattice points, which correspond to all the different $q$-matroids on $\mathbb{F}_{2}^{2}$ and these points are the vertices of $\bar{\mathcal{P}}_{2}^{2}$. In particular, this means that there are no interior lattice points.
Example 3.7.
In this example, we collect some computational results obtained with OSCAR for $\bar{\mathcal{P}}_{2}^{3}$ and $\bar{\mathcal{P}}_{3}^{2}$.
$\bar{\mathcal{P}}_{3}^{2}$ $\bar{\mathcal{P}}_{2}^{3}$
Dimension $5$ $15$
Number of lattice points $7$ $32$
Number of internal lattice points 00 00
Number of vertices $11$ $3483$
We note that also for $\bar{\mathcal{P}}_{3}^{2}$ and $\bar{\mathcal{P}}_{2}^{3}$ we have that all nonzero lattice points are vertices. However, we have, in addition, other rational vertices. This is not just a coincidence, but a property that holds in general. The next result is the main theorem of this section, in which we show that $\mathcal{P}_{q}^{n}$ has no internal lattice points. We denote by $\text{Vert}(\bar{\mathcal{P}}_{q}^{n})$ the set of vertices of $\bar{\mathcal{P}}_{q}^{n}$.
Theorem 3.8.
Let $\mathcal{M}=(\mathcal{L}(E),\rho)$ be a $q$-matroid and let $p_{\mathcal{M}}\in\bar{\mathcal{P}}_{q}^{n}$ be its corresponding point. Then $p_{\mathcal{M}}\in\text{Vert}(\bar{\mathcal{P}}_{q}^{n})$.
Proof.
Let $\mathcal{I}$ and $\mathcal{D}$ denote the collections of independent and dependent spaces of $\mathcal{M}$ respectively. Let $L$ be the loop space of $\mathcal{M}$. By Proposition 3.5 we need to find $T-1$ independent supporting hyperplanes $H_{1},\ldots,H_{T-1}$ such that $\bar{\mathcal{P}}_{q}^{n}\subset H_{i}^{+}$ or $\bar{\mathcal{P}}_{q}^{n}\subset H_{i}^{-}$ for all $i=1,\ldots,T-1$, and such that their intersection is given by the point $p_{\mathcal{M}}$.
So we choose the following hyperplanes:
  1. (1)
    For all $X\in\mathcal{I}\setminus\{\langle 0\rangle\}$ we consider the hyperplane given by the equation $v_{X}=\dim(X)$, having normal vector $e_{X}\in\mathbb{R}^{T-1}$.
  2. (2)
    For all $X\in\mathcal{D}\setminus\mathcal{L}(L)$, let $Y$ be a space in $\mathcal{I}$ such that $Y<X$ and $Y$ is inclusion-maximal in $X$ with this property. We consider then the hyperplane given by $v_{Y}-v_{X}=0$, having normal vector $e_{Y}-e_{X}\in\mathbb{R}^{T-1}$.
  3. (3)
    Finally for all $X\in\mathcal{L}(L)$ we consider the hyperplane given by $v_{X}=0$, having normal vector $e_{X}\in\mathbb{R}^{T-1}$.
Now since $\mathcal{I}\setminus\{\langle 0\rangle\}\cup\mathcal{D}\setminus\mathcal{L}(L)\cup\mathcal{L}(L)=\mathcal{L}(E)^{\ast}$ we have chosen $T-1$ different hyperplanes. Next, we need to show that their normal vectors are linearly independent. Therefore, we write them as rows of a matrix $A$ and, since every $X\in\mathcal{L}(E)^{\ast}$ has a corresponding hyperplane, we can order them, from top to bottom, in the same way we ordered our coordinates (i.e. columns of $A$). Then $A$ is a lower triangular matrix with $1$ and $-1$ on the diagonal. Thus $\det(A)\neq 0$ and so $A$ has rank $T-1$. This implies that the linear system given by $Av=b$ has a unique solution, where the coordinate order of $b$ matches the one of the normal vectors and $b$ has entry $\dim(X)$ if we are in the case (1) and entry 00 anywhere else.
At this point the only thing left to show is that the solution to $Av=b$ is given by our point $p_{\mathcal{M}}$. Using the Gauss-Algorithm for the extended coefficient matrix $(A,b)$ we can eliminate the entries equal to $1$ in rows coming from case (2) in the following way.
Let $R$ such a row, having a $1$ and a $-1$. Then the $1$ corresponds to a space $Y\in\mathcal{I}\setminus\{\langle 0\rangle\}$ s.t. $Y<X$ is inclusion-maximal in $X$, where $X$ corresponds to the $-1$. Since $Y\in\mathcal{I}\setminus\{\langle 0\rangle\}$, there exists a row $\tilde{R}$ coming from the normal vector of a case (1) hyperplane associated to $Y$. Moreover such a row $\tilde{R}$ has $b$ entry equal to $\dim(Y)$. So $R-\tilde{R}$ yields the equation
$-v_{X}=-\dim(Y)\Longleftrightarrow v_{X}=\dim(Y).$
This means that $v_{X}=\rho(X)$ by definition of rank in terms of independent spaces. Now since we chose the hyperplanes of case (2) all in this way, we get $v_{X}=\rho(X)$ for all $X\in\mathcal{D}\setminus\mathcal{L}(L)$. Putting it all together, we get that the solution of $Av=b$ is given by the point $p=(v_{X})_{X\in\mathcal{L}(E)^{\ast}}$ s.t.
$v_{X}=\begin{cases}\;\dim(X)&\text{if }X\in\mathcal{I}\setminus\{\langle 0\rangle\},\\ \;0&\text{if }X\in\mathcal{L}(L),\\ \;\max(\dim(Y)\;|\;Y<X,Y\in\mathcal{I}\setminus\{\langle 0\rangle\})&\text{if }X\in\mathcal{D}\setminus\mathcal{L}(L).\end{cases}$
Thus
$p=(\rho(X))_{X\in\mathcal{L}(E)^{\ast}}=p_{\mathcal{M}}$
and therefore $p_{\mathcal{M}}\in\text{Vert}(\bar{\mathcal{P}}_{q}^{n})$, which completes our proof. ∎
In Theorem 3.8, we showed that $\bar{\mathcal{P}}_{q}^{n}$ has no integer interior points. In particular, the points corresponding to $q$-matroids are necessary vertices in the reduced $q$-rank polytope. We leave as an open question to characterize the rational vertices of $\bar{\mathcal{P}}_{q}^{n}$.
Remark 3.9.
We want to emphasis once more that the $q$-rank polytope $\mathcal{P}_{q}^{n}$ has the same algebraic and geometric properties as the reduced $q$-rank polytope $\bar{\mathcal{P}}_{q}^{n}$. This especially means that Proposition 3.5 and Theorem 3.8 hold for $\mathcal{P}_{q}^{n}$ as well. So, in the remainder of the paper, we apply them also to $\mathcal{P}_{q}^{n}$.
4. The convex combination of two q-(poly)matroids
In this section, we define a new operation on $q$-polymatroids, by looking at their representation as points in the $q$-rank polytope. We initiate the study of the properties of such operation.
Definition 4.1.
Let $\mathcal{M}_{1},\ldots,\mathcal{M}_{t}$ be $t$-many different $q$-polymatroids with ground space $E$. Let $p_{\mathcal{M}_{1}},\ldots,p_{\mathcal{M}_{t}}\in\mathcal{P}_{q}^{n}$ be their corresponding points in the $q$-rank polytope. Finally, let $\lambda_{1},\ldots,\lambda_{t}\in\mathbb{R}_{>0}$ be such that $\sum_{i=1}^{t}\lambda_{i}=1$. A convex combination with coefficients $\lambda_{1},\ldots,\lambda_{t}$ of $\mathcal{M}_{1},\ldots,\mathcal{M}_{t}$ is the $q$-polymatroid $\lambda_{1}\mathcal{M}_{1}+\cdots+\lambda_{t}\mathcal{M}_{t}$, corresponding to the convex combination $\lambda_{1}p_{\mathcal{M}_{1}}+\cdots+\lambda_{t}p_{\mathcal{M}_{2}}\in\mathcal{P}_{q}^{n}$. When we consider a convex combination with coefficients $\lambda_{1},\lambda_{2}$ of two $q$-polymatroid $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, then $\lambda_{2}=1-\lambda_{1}$. Hence, we could fix $\lambda:=\lambda_{1}\in\mathbb{R}$ and talk about a convex combination with coefficient $\lambda$.
Remark 4.2.
Let $\rho_{1},\ldots,\rho_{t}$ be the $q$-rank functions of the $q$-polymatroids $\mathcal{M}_{1},\ldots,\mathcal{M}_{t}$. The $q$-rank function of $\sum\limits_{i=1}^{t}\lambda_{i}\mathcal{M}_{i}$ is given by $\sum_{i=1}^{t}\lambda_{i}\rho_{i}(X)$, for every $X\in\mathcal{L}(E)$. The fact that this is indeed a $q$-rank function is obvious by the definition of $\mathcal{P}_{q}^{n}$.
In the following, we study some structural properties of the $q$-polymatroid arised as convex combination of two given $q$-polymatroids $\mathcal{M}_{1},\mathcal{M}_{2}$. In particular, we investigate the structure of the flats. By restricting to the convex combination of two $q$-matroids, we can also analyze cyclic spaces and $\mu$-independence (in the latter case the coefficients have to be rational).
4.1. The flats
In this subsection, we show that the flats of two $q$-polymatroids are contained in the collection of flats of any of their convex combinations, independently from the choice of coefficients.
Recall that we define a closure operator for $q$-polymatroids as a direct extension of the one for $q$-matroids, as follows.
Definition 4.3.
Let $(\mathcal{L}(E),\rho)$ be a $q$-polymatroid. For each $A\in\mathcal{L}(E)$, define
$\operatorname{Cl}_{\rho}(A):=\{x\in\mathcal{L}(E)\mid\rho(A+x)=r(A)\}.$
The closure function of $(\mathcal{L}(E),\rho)$ is the function defined by
$\operatorname{cl}_{\rho}:\mathcal{L}(E)\to\mathcal{L}(E),\ A\mapsto\operatorname{cl}_{\rho}(A)=\sum_{x\in\operatorname{Cl}_{\rho}(A)}x.$
Let $X\in\mathcal{L}(E)$. It is easy to see that the closure of $X$ is a flat of the $q$-polymatroid $(\mathcal{L}(E),\rho)$, and that if $X$ is a flat, then $X=\operatorname{cl}(X)$. Clearly, $\operatorname{Cl}_{\rho}(F)\supseteq\mathcal{L}(F)_{1}$.
Proposition 4.4.
Let $\mathcal{M}_{1}=(\mathcal{L}(E),\rho_{1}),\;\mathcal{M}_{2}=(\mathcal{L}(E),\rho_{2})$ be two $q$-polymatroids. Let $\mathcal{F}_{i}$ denote the collection of flats of $\mathcal{M}_{i}$, for $i=1,2$. Then, $F\in\mathcal{L}(E)$ is a flat of any convex combination of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ if and only if $\operatorname{Cl}_{\rho_{1}}(F)\cap\operatorname{Cl}_{\rho_{2}}(F)=\mathcal{L}(F)_{1}$.
Proof.
Let $\lambda_{1},\lambda_{2}\in\mathbb{R}_{>0}$, be such that $\lambda_{1}+\lambda_{2}=1$ and $\rho$ be the $q$-rank function of $\lambda_{1}\mathcal{M}_{1}+\lambda_{2}\mathcal{M}_{2}$.
[…]
$\rho(F+x)=\lambda_{1}\rho_{1}(F+x)+\lambda_{2}\rho_{2}(F+x)>\lambda_{1}\rho_{1}(F+x)+\lambda_{2}\rho_{2}(F)\geq\lambda_{1}\rho_{1}(F)+\lambda_{2}\rho_{2}(F).$
A similar reasoning can be done for every $x\in\operatorname{Cl}_{\rho_{2}}(F)$.
($\Longrightarrow$): Assume by contradiction that $\operatorname{Cl}_{\rho_{1}}(F)\cap\operatorname{Cl}_{\rho_{2}}(F)\neq\mathcal{L}(F)_{1}$. Let $x\in\operatorname{Cl}_{\rho_{1}}(F)\cap\operatorname{Cl}_{\rho_{2}}(F)$, then $\rho_{1}(F+x)=\rho_{1}(F)$ and $\rho_{2}(F+x)=\rho_{2}(F)$. Hence
$\rho(F+x)=\lambda_{1}\rho_{1}(F+x)+\lambda_{2}\rho_{2}(F+x)=\lambda_{1}\rho_{1}(F)+\lambda_{2}\rho_{2}(F)=\rho(F),$
from which we conclude that $F$ is not a flat of $\lambda_{1}\mathcal{M}_{1}+\lambda_{2}\mathcal{M}_{2}$. ∎
As a result, we obtain the following immediate corollary.
Corollary 4.5.
Let $\mathcal{M}_{1},\;\mathcal{M}_{2}$ be two $q$-polymatroids. Let $\mathcal{F}_{i}$ denote the collection of flats of $\mathcal{M}_{i}$, for $i=1,2$. Then every $F$ in $\mathcal{F}_{1}\cup\mathcal{F}_{2}$ is a flat of any convex combination of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$.
Proof.
If $F\in\mathcal{F}_{1}\cup\mathcal{F}_{2}$ then $\operatorname{Cl}_{\rho_{1}}(F)\cap\operatorname{Cl}_{\rho_{2}}(F)=\mathcal{L}(F)_{1}$. The rest follows from Proposition 4.4. ∎
4.2. The cyclic spaces
The study of cyclic spaces of the convex combination of two $q$-polymatroids is not as immediate. Hence, in this subsection, we focus on the cyclic spaces of the convex combination of two $q$-matroids.
We introduce the following notation. Let $\mathcal{M}_{1}=(\mathcal{L}(E),\rho_{1}),\mathcal{M}_{2}=(\mathcal{L}(E),\rho_{2})$ be two $q$-matroids. Let $\mathcal{O}_{i}$ denote the collection of cyclic spaces of $\mathcal{M}_{i}$, for $i=1,2$. Further, for a space $X\in\mathcal{L}(E)$ we define $\mathcal{N}_{i}(X)=\{H\in\mathcal{H}(X):\rho_{i}(H)<\rho_{i}(X)\}$, for $i=1,2$.
Proposition 4.6.
Let $\mathcal{M}_{1}=(\mathcal{L}(E),\rho_{1}),\mathcal{M}_{2}=(\mathcal{L}(E),\rho_{2})$ be two $q$-matroids without loops. $X\in\mathcal{L}(E)$ is a cyclic space of any convex combination of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ if and only if $\mathcal{N}_{1}(X)\cap\mathcal{N}_{2}(X)=\emptyset$.
Proof.
Let $\lambda_{1},\lambda_{2}\in\mathbb{R}_{>0}$, be such that $\lambda_{1}+\lambda_{2}=1$ and $\rho$ be the $q$-rank function of $\lambda_{1}\mathcal{M}_{1}+\lambda_{2}\mathcal{M}_{2}$.
[…]
Assume that $X\in\mathcal{O}_{1}\cap\mathcal{O}_{2}$. Then $X$ is a cyclic space in $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, in particular, for every $H\in\mathcal{H}(X)$ we have $\rho_{i}(H)=\rho_{i}(X)$ for $i=1,2$, since $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are $q$-matroids and we have
$\rho(H)=\lambda_{1}\rho_{1}(H)+\lambda_{2}\rho_{2}(H)=\lambda_{1}\rho_{1}(X)+\lambda_{2}\rho_{2}(X)=\rho(X),$
hence $X$ is a cyclic space for $\lambda_{1}\mathcal{M}_{1}+\lambda_{2}\mathcal{M}_{2}$, according to Definition 2.5, part (1).
Now assume that $X\in\mathcal{O}_{1}$ and $X\notin\mathcal{O}_{2}$ (an analogous reasoning can be done for $X\in\mathcal{O}_{2}$ and $X\notin\mathcal{O}_{1}$). Then for every $H\in\mathcal{H}(X)$ we have $\rho_{1}(H)=\rho_{1}(X)$ and there exists an $H^{\prime}\in\mathcal{H}(X)$ such that $\rho_{2}(H^{\prime})<\rho_{2}(X)$. More specifically, since $\mathcal{M}_{2}$ is a $q$-matroid, for every $a\in\mathcal{L}(X)_{1}\setminus\mathcal{L}(H^{\prime})_{1}$ we have $\rho_{2}(X)=\rho_{2}(H^{\prime}+a)=\rho(H^{\prime})+1$. Clearly, part (1) of Definition 2.5 cannot hold for such $H^{\prime}$. We need to verify that for some $a\in\mathcal{L}(X)_{1}\setminus\mathcal{L}(H^{\prime})_{1}$, we have that $0<\rho(X)-\rho(H^{\prime})<\rho(a)$. The first inequality is immediate. For the second one we have
$\rho(X)-\rho(H^{\prime})$ $=\lambda_{1}\rho_{1}(X)+\lambda_{2}\rho_{2}(X)-\lambda_{1}\rho_{1}(H^{\prime})-\lambda_{2}\rho_{2}(H^{\prime})$
$=\lambda_{1}(\rho_{1}(X)-\rho_{1}(H^{\prime}))+\lambda_{2}(\rho_{2}(X)-\rho_{2}(H^{\prime}))$
$=\lambda_{2}<1=\lambda_{1}+\lambda_{2}=\lambda_{1}\rho_{1}(a)+\lambda_{2}\rho_{2}(a)=\rho(a),$
for every $a\in\mathcal{L}(X)_{1}\setminus\mathcal{L}(H^{\prime})_{1}$.
Finally, assume that $X\notin\mathcal{O}_{1}\cup\mathcal{O}_{2}$, but $\mathcal{N}_{1}(X)\cap\mathcal{N}_{2}(X)=\emptyset$. Then, for every $H\in\mathcal{H}(X)$, we have that either $\rho_{1}(H)<\rho_{1}(X)$ or $\rho_{2}(H)<\rho_{2}(X)$, but not both. In particular, for each $H\in\mathcal{H}(X)$ we can argue as in the previous case.
()($\Longrightarrow$)( ⟹ ): Assume by contradiction that there exists some $H^{\prime}\in\mathcal{N}_{1}(X)\cap\mathcal{N}_{2}(X)$. Then $X$ cannot be a cyclic space of $\lambda_{1}\mathcal{M}_{1}+\lambda_{2}\mathcal{M}_{2}$, since for all $a\in\mathcal{L}(X)_{1}\setminus\mathcal{L}(H^{\prime})_{1}$ we would have
$\rho(X)-\rho(H^{\prime})$ $=\lambda_{1}\rho_{1}(X)+\lambda_{2}\rho_{2}(X)-\lambda_{1}\rho_{1}(H^{\prime})-\lambda_{2}\rho_{2}(H^{\prime})$
$=\lambda_{1}(\rho_{1}(X)-\rho_{1}(H^{\prime}))+\lambda_{2}(\rho_{2}(X)-\rho_{2}(H^{\prime}))$
$=\lambda_{1}+\lambda_{2}=\lambda_{1}\rho_{1}(a)+\lambda_{2}\rho_{2}(a)=\rho(a).$
As an immediate result we obtain the following corollary.
Corollary 4.7.
Let $\mathcal{M}_{1},\ \mathcal{M}_{2}$ be two $q$-matroids without loops. Then every $X\in\mathcal{O}_{1}\cup\mathcal{O}_{2}$ is a cyclic space of any convex combination of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$.
Remark 4.8.
If $\mathcal{M}_{1}=(\mathcal{L}(E),\rho_{1})$ or $\mathcal{M}_{2}=(\mathcal{L}(E),\rho_{1})$ have loops we cannot guarantee that all the spaces in $\mathcal{O}_{1}\cup\mathcal{O}_{2}$ are cyclic in a convex combination of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. Let $\lambda_{1},\lambda_{2}\in\mathbb{R}_{>0}$, be such that $\lambda_{1}+\lambda_{2}=1$ and $\rho$ be the $q$-rank function of $\lambda_{1}\mathcal{M}_{1}+\lambda_{2}\mathcal{M}_{2}$. Assume that $\mathcal{M}_{2}$ has no loops, and $X$ is the loop space of $\mathcal{M}_{1}$. Then $X\in\mathcal{O}_{1}$, and assume that $X\notin\mathcal{O}_{2}$. Let $H\in\mathcal{H}(X)$ be such that $\rho_{2}(H)=\rho_{2}(X)-1$. We have that
$\rho(X)-\rho(H^{\prime})$ $=\lambda_{1}\rho_{1}(X)+\lambda_{2}\rho_{2}(X)-\lambda_{1}\rho_{1}(H^{\prime})-\lambda_{2}\rho_{2}(H^{\prime})$
$=\lambda_{1}(\rho_{1}(X)-\rho_{1}(H^{\prime}))+\lambda_{2}(\rho_{2}(X)-\rho_{2}(H^{\prime}))$
$=\lambda_{2}=\lambda_{1}\rho_{1}(a)+\lambda_{2}\rho_{2}(a)=\rho(a),$
for every $a\in\mathcal{L}(X)_{1}\setminus\mathcal{L}(H^{\prime})_{1}$. The last equality is because $\rho_{1}(a)=0$ and $\rho_{2}(a)=1$. In this case, $X$ cannot be a cyclic space for $\lambda_{1}\mathcal{M}_{1}+\lambda_{2}\mathcal{M}2$.
Together the two Corollaries 4.5 and 4.7 immediately yield us the following property of the cyclic flats of such a convex combination.
Corollary 4.9.
Let $\mathcal{M}_{1},\mathcal{M}_{2}$ be two $q$-matroids without loops. Let $\mathcal{Z}_{i}$ denote the collection of cyclic flats of $\mathcal{M}_{i}$, for $i=1,2$. Then every $X\in\mathcal{Z}_{1}\cup\mathcal{Z}_{2}$ is a cyclic flat of any convex combination of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$.
By following [14, Def. 7.7], we introduce the notion of full $q$-polymatroid.
Definition 4.10.
Let $\mathcal{M}$ be a $q$-polymatroid on $E$. We say that $\mathcal{M}$ is full if $\langle 0\rangle$ is a flat and $\mathbb{F}_{q}^{n}$ is a cyclic space of $\mathcal{M}$.
The following result is another direct consequence of Corollary 4.5 and Corollary 4.7.
Proposition 4.11.
Let $\mathcal{M}_{1},\mathcal{M}_{2}$ be two $q$-matroid on E. Assume that $\mathcal{M}_{1}$ is full and $\mathcal{M}_{2}$ has no loops. Then any convex combination $\mathcal{M}$ of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ is full as well.
Proof.
We know that the zero space $\langle 0\rangle$ is a flat of both $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, so by Corollary 4.5 it is a flat of $\mathcal{M}$. Moreover $\mathcal{M}_{1}$ is full, which means in particular that it has no loops and $E$ is a cyclic space of this $q$-matroid. By Corollary 4.7 we then get that $E$ is cyclic in $\mathcal{M}$. ∎
4.3. The $\mu$-independent spaces
In this subsection, we turn to the investigation of the $\mu$-independent spaces of rational $q$-polymatroids arising as convex combination of two $q$-matroids. For this subsection, we let $\mathcal{M}_{1}=(\mathcal{L}(E),\rho_{1}),\mathcal{M}_{2}=(\mathcal{L}(E),\rho_{2})$ be two $q$-matroids. Let $\mathcal{I}_{i}$ denote the collection of independent spaces of $\mathcal{M}_{i}$, for $i=1,2$. Moreover, we set $\mathcal{M}$ to be the convex combination of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ with rational coefficient $10<\lambda<$ 1, as in Definition 4.1. Then $\mathcal{M}$ admits a denominator $\mu$, since $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ correspond to integer points in $\mathcal{P}_{q}^{n}$. Furthermore, we consider the denominator $\mu$ given by the denominator of $\lambda$, which we denote by $\mu=\text{denom}(\lambda)$.
Remark 4.12.
By definition of $\mathcal{M}$ we clearly have that the $\mu=\mathrm{denom}(\lambda)$ always is a denominator for $\mathcal{M}$, but in general such a $\mu$ might not be principal. Nevertheless, Proposition 4.13 shows that there exist infinitely many cases in which the denominator of $\lambda$ is in fact a principal denominator for $\mathcal{M}$.
Proposition 4.13.
If $\mu=\text{denom}(\lambda)$ is prime, then it is principal.
Proof.
Let $\lambda=\frac{a}{b}$, where $b$ is a prime number and $\rho$ be the $q$-rank function of $\mathcal{M}$. By Theorem 3.8 we know that $\mathcal{M}$ corresponds to a rational point in $\mathcal{P}_{q}^{n}$, i.e., there exists a $X\in\mathcal{L}(E)$ such that
$\rho(X)=\lambda\rho_{1}(X)+(1-\lambda)\rho_{2}(X)=\frac{a\rho_{1}(X)+(b-a)\rho_{2}(X)}{b}=\frac{Z}{b}$
and $b$ does not divide $Z$. Moreover $\frac{Z}{b}$ cannot be simplified, since $b$ is a prime number. Let $\mu$ be the principal denominator of $\mathcal{M}$, meaning that $\mu\frac{Z}{b}$ is an integer. So, $b$ has to divide $\mu Z$. Since $b$ is prime, and it does not divide $Z$, it must divide $\mu$. Finally, since $\mu$ is principal and every value of $\rho$ can be written as rational number with denominator $b$, we get that $\mu=b$. ∎
The following proposition describes some of the $\mu$-independent spaces of the $q$-polymatroid $\mathcal{M}$ and it will be useful later on.
Proposition 4.14.
Let $\mathcal{M}_{1}=(\mathcal{L}(E),\rho_{1}),\mathcal{M}_{2}=(\mathcal{L}(E),\rho_{2})$ be two $q$-matroids and $\mathcal{M}$ be their convex combination with coefficient $\lambda$. Let $\mu=\mathrm{denom}(\lambda)$. Then every $X\in\mathcal{I}_{1}\cup\mathcal{I}_{2}$ is a $\mu$-independent space of $\mathcal{M}$, i.e., $\mathcal{I}_{1}\cup\mathcal{I}_{2}\subseteq\mathcal{I}_{\mu}(\mathcal{M})$.
Proof.
Let $\rho$ be the $q$-rank function of $\mathcal{M}$. We distinguish different cases. First assume that $X\in\mathcal{I}_{1}\cap\mathcal{I}_{2}$, then we know that $\rho_{1}(Y)=\rho_{2}(Y)=\dim(Y)$ for every $Y\leq X$. Thus
$\rho(Y)=\lambda\rho_{1}(Y)+(1-\lambda)\rho_{2}(Y)=\dim(Y),$
and we get $\rho(Y)\geq\frac{\dim(Y)}{\mu}$. Therefore $X$ is $\mu$-independent. Next assume that $X\in\mathcal{I}_{1}$ but $X\not\in\mathcal{I}_{2}$. This means that we have $\rho_{1}(Y)=\dim(Y)$ for all $Y\leq X$. Now since $\rho_{2}(X)>0$ we compute
$\rho(Y)=\lambda\rho_{1}(Y)+(1-\lambda)\rho_{2}(Y)>\lambda\rho_{1}(Y)=\lambda\dim(Y)\geq\frac{\dim(Y)}{\mu},$
where the last inequality follows from the fact that $\lambda=\frac{a}{\mu}$, for some $a\geq 1$. Thus, also in this case, $X$ is $\mu$-independent. Finally, for the case $X\in\mathcal{I}_{2}$ but $X\not\in\mathcal{I}_{1}$ we can argue in the same way as above. ∎
Remark 4.15.
Note that there exist $\mu$-independent spaces in $\mathcal{M}$ which are neither independent in $\mathcal{M}_{1}$ nor in $\mathcal{M}_{2}$. However, these spaces are not easy to classify, since the $\mu$-independence in these cases highly depends on the denominator $\mu$, the ambient dimension $n$ and the ranks of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. We refer to Example 6.3, for seeing this property.
The next corollary is a consequence of Proposition 4.14 and characterizes some properties of $\mu$-dependent spaces and $\mu$-circuits. We fix the notation $\mathcal{D}_{i}$ for the collection of dependent spaces and $\mathcal{C}_{i}$ for the collection of circuits of $\mathcal{M}_{i}$, for $i=1,2$.
Corollary 4.16.
Let $\mathcal{D}_{\mu}(\mathcal{M})$ and $\mathcal{C}_{\mu}(\mathcal{M})$ be respectively the collections $\mu$-dependent spaces and $\mu$-circuits of $\mathcal{M}$. Then it holds that
  1. (1)
    $\mathcal{D}_{\mu}(\mathcal{M})\subseteq\mathcal{D}_{1}\cap\mathcal{D}_{2}$ and
  2. (2)
    $(\mathcal{C}_{1}\cup\mathcal{C}_{2})\cap\mathcal{D}_{\mu}\subseteq\mathcal{C}_{\mu}(\mathcal{M})$.
Proof.
The first statement follows by taking the complement of both sides of the expression $\mathcal{I}_{1}\cup\mathcal{I}_{2}\subseteq\mathcal{I}_{\mu}(\mathcal{M})$ from Proposition 4.14. For the second statement we have to consider three cases. Assume first that that $X\in(\mathcal{C}_{1}\cap\mathcal{C}_{2})\cap\mathcal{D}_{\mu}$. So we know that $X$ is $\mu$-dependent and all its proper subspace $Y<X$ are contained in $\mathcal{I}_{1}\cap\mathcal{I}_{2}$, thus they are all $\mu$-independent by Proposition 4.14. This implies that $X$ is a $\mu$-circuit. Now we assume that $X\in\mathcal{C}_{1}\cap\mathcal{D}_{\mu}$ but $X\not\in\mathcal{C}_{2}$. Once more we get that $X$ is $\mu$-dependent, moreover we know that all its proper subspaces $Y<X$ are contained in $\mathcal{I}_{1}$. Then by using Proposition 4.14 they are $\mu$-independent and so $X$ is a $\mu$-circuit. Lastly, for the case $X\in\mathcal{C}_{2}\cap\mathcal{D}_{\mu}$ but $X\not\in\mathcal{C}_{1}$, we can argue as above. ∎
A more precise characterization of the collections $\mathcal{D}_{\mu}(\mathcal{M})$ and $\mathcal{C}_{\mu}(\mathcal{M})$ would be more involved due to the issues addressed in Remark 4.15.
5. The convex combination of paving q-matroids
In this section we focus on the convex combination of special $q$-matroids, namely those that are paving. We recall that a $q$-matroid $\mathcal{M}$ is said to be paving when any of its circuits $C$ satisfies $\dim(C)\geq\text{rank}(\mathcal{M})$; see [13]. We generalize the pavingness property for rational $q$-polymatroids. Afterwards, we show that any convex combination of paving $q$-matroids with the same ranks gives rise to paving $q$-polymatroids. Furthermore, we discuss the flats and cyclic spaces of such a convex combination.
We start with an explicit construction of paving $q$-matroids, which was proven in [12].
Proposition 5.1.
[12, Prop 4.6.] Let $11\leq k\leq n-$ 1. In addition, let $\mathcal{S}\subseteq\mathcal{L}(E)_{k}$, such that $\dim(V\cap W)\leq k-2$, for every two distinct $V,W\in\mathcal{S}$. Define the map
$\rho_{\mathcal{S}}:\mathcal{L}(E)\rightarrow\mathbb{Z}_{\geq 0},\quad V\mapsto\begin{cases}\;\hfil k-1\;&\text{if }V\in\mathcal{S},\\ \;\hfil\min\{\dim V,k\}\;&\text{otherwise}.\end{cases}$
Then $\mathcal{M}_{\mathcal{S}}=(\mathcal{L}(E),\rho_{\mathcal{S}})$ is a paving $q$-matroid of rank $k$, whose circuits of rank $k-1$ are the subspaces in $\mathcal{S}$. We call $\mathcal{M}_{\mathcal{S}}$ the paving $q$-matroid induced by $\mathcal{S}$.
Remark 5.2.
Not all paving $q$-matroids arise from Proposition 5.1. We will restrict our study only to this class, since we know exactly the “shape” of their corresponding points in the $q$-rank polytope $\mathcal{P}_{q}^{n}$.
A special example of paving $q$-matroids constructed as in Proposition 5.1 is given below.
Example 5.3.
Let $\mathcal{S}=\emptyset\subset\mathcal{L}(E)_{k}$, for some fix $11\leq k\leq n-$ 1. Then the $q$-rank function from Proposition 5.1 simplifies to $\rho(V)=\min\{\dim V,k\}$, for all $V\in\mathcal{L}(E)$. This is precisely the uniform $q$-matroid $\mathcal{U}_{k,n}(q)$, having all the $(k+1)$-dimensional space of $\mathbb{F}_{q}^{n}$ as circuits. So, we can regard the uniform $q$-matroid $\mathcal{U}_{k,n}(q)$ as $\mathcal{M}_{\emptyset}$, the paving $q$-matroid induced by $\emptyset$.
The next definition introduces the notion of pavingness for the more general case of rational $q$-polymatroids.
Definition 5.4.
Let $\mathcal{M}$ be a rational $q$-polymatroid on $E$ with denominator $\mu$. Denote by $\mathcal{I}_{\mu}(\mathcal{M})$ its collection of $\mu$-independent spaces. Then we say that $\mathcal{M}$ is $\mu$-paving if for all $\mu$-circuits $C\in\mathcal{L}(E)$ it holds that $\dim(C)\geq\max\{\dim(V)\;|\;V\in\mathcal{I}_{\mu}(\mathcal{M})\}$.
Remark 5.5.
We want to emphasis that this notion of pavingness is not a simple generalization of the same definition for $q$-matroids. The reason is that the rank of a $q$-polymatroid is given by the rank of its $\mu$-bases, which are the maximal dimensional $\mu$-independent spaces. However, this rank is not necessarily equal to their dimension, which implies
$\max\{\dim(V)\;|\;V\in\mathcal{I}_{\mu}(\mathcal{M})\}\geq\text{rank}(\mathcal{M}).$
Therefore, if a $q$-polymatroid is paving according to Definition 5.4, then the dimension of every $\mu$-circuit is greater or equal than the rank of the $q$-polymatroid. A detailed discussion about $\mu$-independent spaces can be found in [11, Sec. 3-4].
For the remainder of the section, we fix the following setting. Let $11\leq k\leq n-$ 1 and $\mathcal{S}_{1},\mathcal{S}_{2}\subseteq\mathcal{L}(E)_{k}$, with the property described in Proposition 5.1. Assume for simplicity that $\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset$ (we address the other case later on). For $i=1,2$ let $\mathcal{M}_{i}=\mathcal{M}_{\mathcal{S}_{i}}$ be the paving $q$-matroid induced by $\mathcal{S}_{i}$ and $p_{\mathcal{M}_{i}}$ be its corresponding point in $\mathcal{P}_{q}^{n}$. Moreover set $\mathcal{M}$ to be the convex combination of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ with rational coefficient $10<\lambda<$ 1, as in Definition 4.1 and again consider the denominator $\mu=\text{denom}(\lambda)$ for $\mathcal{M}$. Finally denote by $p_{\mathcal{M}}\in\mathcal{P}_{q}^{n}$ the corresponding point of $\mathcal{M}$.
The following theorem shows that the pavingness property is preserved under taking convex combinations.
Theorem 5.6.
Let $\mathcal{M}_{1},\mathcal{M}_{2}$ and $\mathcal{M}$ as defined in the above setting. Then $\mathcal{M}$ is $\mu$-paving.
Before proving Theorem 5.6 we first collect some useful properties concerning the shape of the points $p_{\mathcal{M}_{1}},\;p_{\mathcal{M}_{2}}$ and $p_{\mathcal{M}}$.
Remark 5.7.
By Proposition 5.1, we immediately get that for $i=1,2$,
$p_{\mathcal{M}_{i}}=(p_{\mathcal{M}_{i},X})_{X\in\mathcal{L}(E)}=\begin{cases}\;\hfil k-1\;&\text{if }X\in\mathcal{S}_{i},\\ \;\hfil\dim(X)\;&\text{if }\dim(X)\leq k,\;X\not\in\mathcal{S}_{i},\\ \;\hfil k\;&\text{otherwise}.\end{cases}$
Therefore the point $p_{\mathcal{M}}$ corresponding to the convex combination of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ is given by
$p_{\mathcal{M}}=(p_{\mathcal{M},X})_{X\in\mathcal{L}(E)}=\begin{cases}\;\hfil\dim(X)\;&\text{if }\dim(X)\leq k,\;X\not\in\mathcal{S}_{1}\cup\mathcal{S}_{2},\\ \;\hfil\lambda(k-1)+(1-\lambda)k\;&\text{if }X\in\mathcal{S}_{1},\\ \;\hfil\lambda k+(1-\lambda)(k-1)\;&\text{if }X\in\mathcal{S}_{2},\\ \;\hfil k\;&\text{if }\dim(X)\geq k+1.\end{cases}$
Proof of Theorem 5.6.
Let $\mathcal{I}_{1},\mathcal{I}_{2}$ be the collections of independent spaces of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ respectively and set $\mathcal{I}_{\mu}(\mathcal{M})$ to be the collection of $\mu$-independent spaces of $\mathcal{M}$. By Remark 5.7 we know that every space $X\not\in\mathcal{S}_{i}$ of dimension $\dim(X)\leq k$ is independent in $\mathcal{M}_{i}$, for $i=1,2$. Thus we get that $\mathcal{I}_{1}\cup\mathcal{I}_{2}=\mathcal{L}(E)_{\leq k}$, since $\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset$. Furthermore, that all spaces $X$, having dimension $\dim(X)\leq k$, are $\mu$-independent spaces of $\mathcal{M}$, by Proposition 4.14. Now, the spaces $X\in\mathcal{L}(E)_{s}$ with $s\in\{k+1,\ldots,n\}$ have the same rank value, namely $p_{\mathcal{M},X}=k$, from Remark 5.7. In particular, for a fixed such $s$, either every $X\in\mathcal{L}(E)_{s}$ satisfies
$p_{\mathcal{M},X}\geq\frac{\dim(X)}{\mu},$
or none of them does. In the first case clearly every subspace $Y$ s.t. $\dim(Y)\in\{k+1,\ldots,s-1\}$ would then also satisfies this inequality and therefore we would get
$\bigcup_{t=k+1}^{s}\mathcal{L}(E)_{s}\subseteq\mathcal{I}_{\mu}(\mathcal{M}).$
In the second case, we would declare all of those spaces $X$ to be $\mu$-dependent. All together this implies that the $\mu$-circuits are given by the collection of all $s_{0}$-dimensional spaces such that $s_{0}\in\{k+1,\ldots,n\}$ is the smallest dimension for which the inequality gets violated the first time and moreover that
$\mathcal{I}_{\mu}(\mathcal{M})=\bigcup_{t=0}^{s_{0}-1}\mathcal{L}(E)_{t}.$
Thus, the maximal dimension of any $\mu$-independent space is $s_{0}-1<s_{0}$ and so $\mathcal{M}$ is $\mu$-paving according to Definition 5.4. ∎
As a byproduct, we obtain the collections of the $\mu$-independent spaces and $\mu$-circuits of $\mathcal{M}$.
Corollary 5.8.
Let $\mathcal{M}_{1},\mathcal{M}_{2}$ and $\mathcal{M}$ as before. Let $s_{0}\in\{k+1,\ldots,n\}$ be the smallest integer such that there exists a space $X\in\mathcal{L}(E)_{s_{0}}$ violating the inequality $p_{\mathcal{M},X}\geq\dim(X)/\mu$. Then the following holds:
$\mathcal{I}_{\mu}(\mathcal{M})=\mathcal{L}(E)_{\leq s_{0}-1}\quad\text{ and }\quad\mathcal{C}_{\mu}(\mathcal{M})=\mathcal{L}(E)_{s_{0}}.$
Remark 5.9.
Theorem 5.6 still holds if we replace the assumption $\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset$ with the weaker one $\mathcal{S}_{1}\not=\mathcal{S}_{2}$. The reason is that the additional possibility $p_{\mathcal{M},X}=k-1$ for $X\in\mathcal{S}_{1}\cap\mathcal{S}_{2}$ also yields that $X$ is $\mu$-independent. This is due the fact that $\mu\geq 2$, while, at the same time, $\frac{k}{k-1}<2$ and consequently we have $k-1\geq k/\mu$. Therefore, all space of dimension smaller or equal than $k$ are $\mu$-independent in $\mathcal{M}$ and we arrive again at the starting point of the proof of Theorem 5.6.
For the remainder of this subsection we focus on the characterization of the flats, cyclic spaces and cyclic flats of such a convex combination $\mathcal{M}$ arising from two paving $q$-matroids $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. We start by describing these collections for general $q$-matroids coming from the construction in Proposition 5.1.
Lemma 5.10.
Let $11\leq k\leq n-$ 1 and $\mathcal{S}\subseteq\mathcal{L}(E)_{k}$ with the property described in Proposition 5.1. Moreover, let $\mathcal{M}_{S}$ be the paving $q$-matroid induced by $\mathcal{S}$ and denote by $\mathcal{F}_{\mathcal{M}_{S}},\mathcal{O}_{\mathcal{M}_{S}}$ and $\mathcal{Z}_{\mathcal{M}_{S}}$ its collections of flats, cyclic spaces and cyclic flats respectively. Then the following holds:
  1. (1)
    The flats of $\mathcal{M}_{S}$ are given by
    $\mathcal{F}_{\mathcal{M}_{\mathcal{S}}}=(\mathcal{L}(E)_{\leq k-1}\setminus I)\cup S\cup\{\mathbb{F}_{q}^{n}\},$
    where $I$ is given by
    $I:=\{H\in\mathcal{L}(E)_{k-1}\;|\;H\text{ is contained in some }S\in\mathcal{S}\}.$
  2. (2)
    The cyclic spaces of $\mathcal{M}_{S}$ are given by
    $\mathcal{O}_{\mathcal{M}_{\mathcal{S}}}=(\mathcal{L}(E)_{\geq k+1}\setminus J)\cup S\cup\{\langle 0\rangle\},$
    where $J$ is given by
    $J:=\{Y\in\mathcal{L}(E)_{k+1}\;|\;Y\text{ contains some }S\in\mathcal{S}\}.$
  3. (3)
    The cyclic flats are given by
    $\mathcal{Z}_{\mathcal{M}_{\mathcal{S}}}=\mathcal{S}\cup\{E\}\cup\{\langle 0\rangle\}.$
Proof.
(1): By Remark 5.7, we have that every space $X\in\mathcal{L}(E)_{\geq k+1}$ has the same rank, so they cannot be flats of $\mathcal{M}_{\mathcal{S}}$, expcept for the whole space $E$. Next we have that all $k$-dimensional spaces are either bases of $\mathcal{M}_{\mathcal{S}}$ or members of $\mathcal{S}$. While bases can never be flats, all the members of $\mathcal{S}$ are flats, since their rank increases from $k-1$ to $k$ whenever we add any $1$-dimensional space not already contained in them. Let $X\in\mathcal{L}(E)_{k-1}$. Then, $X$ is independent. If $X$ is not contained in any element of $\mathcal{S}$, then it is clearly a flat of $\mathcal{M}_{\mathcal{S}}$. If $X$ is an hyperplane of some element of $\mathcal{S}$, we observe that there exists a $1$-dimensional space that added to $X$ gives a space of the same rank as $X$. Hence, such spaces cannot be flats. Finally all the spaces $X\in\mathcal{L}(E)_{\leq k-2}$ are independent and so are all the $(\dim(X)+1)$-dimensional spaces containing them, thus they are flats of $\mathcal{M}_{\mathcal{S}}$.
(2): Cyclic spaces cannot be independent, except from the zero-space. So, we only need to consider dependent spaces. Clearly, circuits are cyclic, thus all elements of $\mathcal{S}$ are cyclic in $\mathcal{M}_{\mathcal{S}}$. Let $X\in\mathcal{L}(E)_{k+1}$. We distinguish two cases. If $X$ does not contain any element from $\mathcal{S}$, then it is cyclic, by Remark 5.7. Assume that $X$ contains an element from $S\in\mathcal{S}$. Then $S$ has rank strictly smaller that the rank of $X$, by Remark 5.7, and therefore $X$ cannot be cyclic in $\mathcal{M}_{\mathcal{S}}$. Finally, all spaces $X\in\mathcal{L}(E)_{\geq k+2}$ have rank $k$ and so do all their hyperplanes, which implies that they are cyclic in $\mathcal{M}_{\mathcal{S}}$. All together this proves the second statement.
(3): This statement follows immediately by intersecting the sets $\mathcal{F}_{\mathcal{M}_{\mathcal{S}}}$ and $\mathcal{O}_{\mathcal{M}_{\mathcal{S}}}$. ∎
With the aid of Lemma 5.10 we can characterize the flats, cyclic spaces and cyclic flats of the convex combination of two paving $q$-matroids. Therefore let $\mathcal{M}_{1},\mathcal{M}_{2}$ and $\mathcal{M}$ be as before.
Theorem 5.11.
Let $\mathcal{F}_{i}$ denote the collection of flats of $\mathcal{M}_{i}$, for $i=1,2$. Then the collection of flats of $\mathcal{M}$ is given by
$\mathcal{F}_{\mathcal{M}}=\mathcal{S}_{1}\cup\mathcal{S}_{2}\cup\{E\}\cup\mathcal{L}(E)_{\leq k-1}.$
Proof.
For any $X\in\mathcal{L}(E)$ let $\text{Cl}_{1}(X)$ and $\text{Cl}_{2}(X)$ the sets defined in Definition 4.3 corresponding to $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, respectively. By Corollary 4.5, we already have $\mathcal{F}_{1}\cup\mathcal{F}_{2}\subseteq\mathcal{F}_{\mathcal{M}}$. Using Lemma 5.10 (1) for $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ individually one computes the following:
$\mathcal{F}_{1}\cup\mathcal{F}_{2}$ $=\mathcal{S}_{1}\cup\mathcal{S}_{2}\cup\{E\}\cup(\mathcal{L}(E)_{\leq k-1}\setminus I_{1})\cup(\mathcal{L}(E)_{\leq k-1}\setminus I_{2})$
$=\mathcal{S}_{1}\cup\mathcal{S}_{2}\cup\{E\}\cup(\mathcal{L}(E)_{\leq k-1}\setminus(I_{1}\cap I_{2})).$
Here $I_{1}$ and $I_{2}$ denote the corresponding sets from Lemma 5.10(1). For all the other spaces $F\in\mathcal{L}(E)\setminus(\mathcal{F}_{1}\cup\mathcal{F}_{2})$ we still need to check if they are flats of $\mathcal{M}$. Due to Proposition 4.4, this just means to check if $\text{Cl}_{1}(F)\cap\text{Cl}_{2}(F)=\mathcal{L}(F)_{1}$ holds. There exist three cases we have to consider:
  • $F\in\mathcal{L}(E)_{k}\setminus(\mathcal{S}_{1}\cup\mathcal{S}_{2})$: This implies that $F$ is basis in both $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ and consequently we get $\text{Cl}_{1}(F)=\text{Cl}_{2}(F)=\mathcal{L}(E)_{1}$. Thus $F$ is not a flat of $\mathcal{M}$.
  • $F\in\mathcal{L}(E)_{\geq k+1}\setminus\{\mathbb{F}_{q}^{n}\}$: In this case we know from Remark 5.7 that $p_{\mathcal{M}_{1},F}=p_{\mathcal{M}_{2},F}=k$ and this is also the rank of all the spaces containing it. Hence, $\text{Cl}_{1}(F)=\text{Cl}_{2}(F)=\mathcal{L}(E)_{1}$ and $F$ is not a flat of $\mathcal{M}$.
  • $F\in I_{1}\cap I_{2}$: This means that there exist spaces $X_{1}\in\mathcal{S}_{1}$ and $X_{2}\in\mathcal{S}_{2}$ s.t. $F\leq X_{1}\cap X_{2}$. Moreover $X_{1}\not=X_{2}$ since $\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset$ by assumption. Now if there exists a $1$-dimensional space $x\in\mathcal{L}(E)_{1}\setminus\mathcal{L}(F)_{1}$ s.t. $F+x=X_{1}$ and $F+x=X_{2}$, we get that $X_{1}=X_{2}$, which contradicts the assumption. Therefore none of the $1$-dimensional spaces $x\in\mathcal{L}(E)_{1}\setminus\mathcal{L}(F)_{1}$ can be in $\text{Cl}_{1}(F)\cap\text{Cl}_{2}(F)$, implying that $\text{Cl}_{1}(F)\cap\text{Cl}_{2}(F)=\mathcal{L}(F)_{1}$ and hence $F$ is a flat of $\mathcal{M}$.
By combining these three cases with the above computation of $\mathcal{F}_{1}\cup\mathcal{F}_{2}$ we then finally obtain the desired expression. ∎
Theorem 5.12.
Let $\mathcal{O}_{i}$ denote the collection of cyclic spaces of $\mathcal{M}_{i}$, for $i=1,2$. Then the collection of cyclic spaces of $\mathcal{M}$ is given by
$\mathcal{O}_{\mathcal{M}}=\mathcal{S}_{1}\cup\mathcal{S}_{2}\cup\{\langle 0\rangle\}\cup\mathcal{L}(E)_{\geq k+1}.$
Proof.
First of all notice that $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ do not have loops, since they are paving. Therefore we can apply Corollary 4.7, which yields $\mathcal{O}_{1}\cup\mathcal{O}_{2}\subseteq\mathcal{O}_{\mathcal{M}}$. Then, using Lemma 5.10 (2) for $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ individually one computes the following:
$\mathcal{O}_{1}\cup\mathcal{O}_{2}$ $=\mathcal{S}_{1}\cup\mathcal{S}_{2}\cup\{\langle 0\rangle\}\cup(\mathcal{L}(E)_{\geq k+1}\setminus J_{1})\cup(\mathcal{L}(E)_{\geq k+1}\setminus J_{2})$
$=\mathcal{S}_{1}\cup\mathcal{S}_{2}\cup\{\langle 0\rangle\}\cup(\mathcal{L}(E)_{\geq k+1}\setminus(J_{1}\cap J_{2})).$
Here $J_{1}$ and $J_{2}$ denote the corresponding sets from Lemma 5.10 (2). For all the other spaces $V\in\mathcal{L}(E)\setminus(\mathcal{O}_{1}\cup\mathcal{O}_{2})$ we still need to check if they are cyclic spaces of $\mathcal{M}$. Due to Proposition 4.6, this just means to check if $\mathcal{N}_{1}(V)\cap\mathcal{N}_{2}(V)=\emptyset$ holds, where $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ are the sets defined in Section 4.2. We need to distinguish between three cases:
  • $V\in\mathcal{L}(E)_{k}\setminus(\mathcal{S}_{1}\cup\mathcal{S}_{2})$: This implies that $V$ is basis in both $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ and consequently we get $\mathcal{N}_{1}(V)=\mathcal{N}_{2}(V)=\mathcal{H}(V)$. Thus $V$ is not a cyclic space of $\mathcal{M}$.
  • $V\in\mathcal{L}(E)_{\leq k-1}\setminus\{\langle 0\rangle\}$: In this case we know that $V$ is independent in $\mathcal{M}_{1}$ as well as in $\mathcal{M}_{2}$, hence we have again $\mathcal{N}_{1}(V)=\mathcal{N}_{2}(V)=\mathcal{H}(V)$ and $V$ is not a cyclic space of $\mathcal{M}$.
  • $V\in J_{1}\cap J_{2}$: This means that $V$ has some hyperplanes contained in $\mathcal{S}_{1}$ and also some contained $\mathcal{S}_{2}$. Moreover by the definition of $\mathcal{N}_{i}$ we can observe that $\mathcal{N}_{i}(V)\subseteq\mathcal{S}_{i}$, for $i=1,2$. Thus there exists no hyperplane contained in $\mathcal{N}_{1}(V)\cap\mathcal{N}_{2}(V)$, since $\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset$. In other words we have $\mathcal{N}_{1}(V)\cap\mathcal{N}_{2}(V)=\emptyset$ and $V$ is cyclic space of $\mathcal{M}$.
By combining these three cases with the above computation of $\mathcal{O}_{1}\cup\mathcal{O}_{2}$ we then finally obtain the desired expression. ∎
As a direct consequence of the above theorems we get the following description of the cyclic flats of $\mathcal{M}$.
Corollary 5.13.
Let $\mathcal{Z}_{i}$ denote the collection of cyclic flats of $\mathcal{M}_{i}$, for $i=1,2$. Then the collection of cyclic flats of $\mathcal{M}$ is given by
$\mathcal{Z}_{\mathcal{M}}=\mathcal{S}_{1}\cup\mathcal{S}_{2}\cup\{\langle 0\rangle\}\cup\{E\}.$
Proof.
Intersecting the sets $\mathcal{F}_{\mathcal{M}}$ and $\mathcal{O}_{\mathcal{M}}$ obtained by Theorem 5.11 and Theorem 5.12 respectively yields the claim. ∎
Note that by weakening the assumption $\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset$ in Theorem 5.11, Theorem 5.12 and Corollary 5.13 to the condition $\mathcal{S}_{1}\not=\mathcal{S}_{2}$, the computation of these collections becomes more complicated and the results will not have such a compact form anymore.
We end this subsection with one final remark on this kind of the convex combinations.
Remark 5.14.
The reader may have noticed that our setting only considers two paving $q$-matroids coming from the construction in Proposition 5.1, which in addition have the same rank. One can also do it for two such paving $q$-matroids of arbitrary rank. However, all the computations become significantly more difficult, as the shape of the points $p_{\mathcal{M}_{1}},p_{\mathcal{M}_{2}}$ and $p_{\mathcal{M}}$ include several new cases to consider. For this reason we decided to not discuss this situation in general. However, for uniform $q$-matroids we obtain interesting results, see Section 6.
6. The convex combination of uniform q-matroids
In this section, we first describe the convex combination with rational coefficients of two uniform $q$-matroids of any rank, by investigating its $\mu$-independent spaces. Then, we turn to the convex combination of $(n-2)$ many uniform $q$-matroids with rank strictly increasing. Finally, we characterize the cyclic flats of such convex combinations.
Recall that if $\mathcal{M}$ is the uniform $q$-matroid $\mathcal{U}_{k,n}(q)$, then its corresponding point $p_{\mathcal{M}}\in\mathcal{P}_{q}^{n}$ is
$p_{\mathcal{M}}=(0,1,\ldots,1,2,\ldots,2,\ldots,k,\ldots,k,\dots,k),$ (1)
where the first coordinate is a 00, the successive $\left[\begin{matrix}n\\ 1\end{matrix}\right]_{q}$ coordinates are equal to $1$, and so on, till the last $\sum\limits_{i=k}^{n}\left[\begin{matrix}n\\ i\end{matrix}\right]_{q}$ positions are equal to $k$.
6.1. Convex combination of two uniform q-matroids
In this subsection, we consider two uniform $q$-matroids. Let $1<k_{1}<k_{2}<n$ and $\mathcal{M}_{1}=\mathcal{U}_{k_{1},n}(q)$, $\mathcal{M}_{2}=\mathcal{U}_{k_{2},n}(q)$. Let $p_{\mathcal{M}_{1}}$ and $p_{\mathcal{M}_{2}}$ be of the form described in Eq. (1). Let $\lambda=\frac{a}{\mu}\in\mathbb{Q}$, with $a\geq 1$ and $\mu\geq 2$ Consider the convex combination $\mathcal{M}=(1-\lambda)\mathcal{M}_{1}+\lambda\mathcal{M}_{2}$ and its corresponding point $p_{\mathcal{M}}\in\mathcal{P}_{q}^{n}$. We have already seen that $\mu$ is a denominator for $\mathcal{M}$. Hence, we can study the $\mu$-independence.
Theorem 6.1.
Let $a,\mu$ be integers, with $a\geq 1,\;\mu\geq 2$ and $\lambda=\frac{a}{\mu}\in\mathbb{Q}$ and $\mathcal{M}=(1-\lambda)\mathcal{M}_{1}+\lambda\mathcal{M}_{2}$. If $\mu\geq\left\lceil\frac{n}{k_{1}}\right\rceil$, then $\mathcal{I}_{\mu}(\mathcal{M})=\mathcal{L}(E)$.
Proof.
Let $p_{\mathcal{M}}=(p_{X})_{X\in\mathcal{L}(E)}$ be the point in $\mathcal{P}_{q}^{n}$ corresponding to $\mathcal{M}$. Then
$p_{X}=\begin{cases}\dim(X)&\textnormal{ if }\dim(X)\leq k_{1},\\ k_{1}+\lambda(\dim(X)-k_{1})&\textnormal{ if }k_{1}<\dim(X)\leq k_{2},\\ (1-\lambda)k_{1}+\lambda k_{2}&\textnormal{ if }\dim(X)>k_{2}.\end{cases}$
From Proposition 4.14, we have that the independent spaces in $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are $\mu$-independent spaces in $\mathcal{M}$. In particular, all the spaces $X\in\mathcal{L}(E)_{\leq k_{2}}$ are $\mu$-independent. Let $X\in\mathcal{L}(E)_{>k_{2}}$. Then $p_{X}=k_{1}+\lambda(k_{2}-k_{1})$. We need to show that for every subspace $J\leq X$, we have that $p_{J}\geq\frac{\dim(J)}{\mu}$. This is clearly true for subspaces of dimension at most $k_{2}$. So, we consider $J\leq X$ with $\dim(J)>k_{2}$. Then
$p_{J}$ $=(1-\lambda)k_{1}+\lambda k_{2}=k_{1}+\lambda(k_{2}-k_{1})=k_{1}+\frac{a}{\mu}(k_{2}-k_{1})$
$\geq k_{1}+\frac{k_{2}-k_{1}}{\mu}=\frac{(\mu-1)k_{1}+k_{2}}{\mu},$
where the inequality follows from the fact that $a\geq 1$. Now, we have that
$\frac{(\mu-1)k_{1}+k_{2}}{\mu}\geq\frac{\dim(J)}{\mu}$ $\Leftrightarrow(\mu-1)k_{1}+k_{2}\geq\dim(J)\Leftrightarrow\frac{\dim(J)-k_{2}}{k_{1}}+1\leq\mu.$ (2)
Observe that, since $k_{1}<k_{2}$, we get
$\frac{\dim(J)-k_{2}}{k_{1}}+1\leq\frac{\dim(J)-k_{1}}{k_{1}}+1=\frac{\dim(J)}{k_{1}}\leq\left\lceil\frac{n}{k_{1}}\right\rceil\leq\mu,$ (3)
where the last inequality follows by assumption. Hence, by putting together Eq. (2) and Eq. (3), we conclude that $X$ is $\mu$-independent. ∎
The assumption $\mu\geq\left\lceil\frac{n}{k_{1}}\right\rceil$ in Theorem 6.1 is quite strong, but there are some cases where we can show the same result for every choice of $\lambda$, as the following theorem illustrates.
Theorem 6.2.
Let $1<k_{1}<k_{2}<n$, with $k_{1}+k_{2}\geq n$ and let $\mathcal{M}_{1}=\mathcal{U}_{k_{1},n}(q)$, $\mathcal{M}_{2}=\mathcal{U}_{k_{2},n}(q)$. Let $\lambda=\frac{a}{\mu}\in\mathbb{Q}$ and $\mathcal{M}=(1-\lambda)\mathcal{M}_{1}+\lambda\mathcal{M}_{2}$. then $\mathcal{I}_{\mu}(\mathcal{M})=\mathcal{L}(E)$.
Proof.
Let $p_{\mathcal{M}}=(p_{X})_{X\in\mathcal{L}(E)}$ be the point in $\mathcal{P}_{q}^{n}$ corresponding to $\mathcal{M}$. As in the proof of Theorem 6.1, it is enough the show that spaces of dimension larger than $k_{2}$ are $\mu$-independent. Let $X\in\mathcal{L}(E)_{>k_{2}}$. Then $p_{X}=k_{1}+\lambda(k_{2}-k_{1})$. We need to show that for every subspace $J\leq X$, we have that $p_{J}\geq\frac{\dim(J)}{\mu}$. This is clearly true for subspaces of dimension at most $k_{2}$. So, we consider $J\leq X$ with $\dim(J)>k_{2}$. Then
$p_{J}$ $=\left(1-\frac{a}{\mu}\right)k_{1}+\frac{a}{\mu}k_{2}=\frac{\mu-a}{\mu}k_{1}+\frac{a}{\mu}k_{2}.$
Observe that since $a\geq 1$ and $\mu\geq 2$,
$(\mu-a)k_{1}+ak_{2}=\mu k_{1}+a(k_{2}-k_{1})\geq 2k_{1}+k_{2}-k_{1}=k_{1}+k_{2}\geq n\geq\dim(J).$
By dividing by $\mu$ both sides of the inequality, we have that $X$ is $\mu$-independent. ∎
Note that while the assumptions of Theorem 6.1 and Theorem 6.2 are sufficient to characterize the $\mu$-independent spaces of the convex combination of two uniform $q$-matroids, they are not necessary, as next example shows.
Example 6.3.
Let $k_{1}=2,\;k_{2}=3,\;n=8$ and $\mathcal{M}_{1}=\mathcal{U}_{2,8}(q)$, $\mathcal{M}_{1}=\mathcal{U}_{3,8}(q)$. Let $\lambda=\frac{2}{3}$. Then $\mu=3<4=\frac{n}{k_{1}}$. Moreover, for every subspace $X$ of $\mathbb{F}_{q}^{8}$ with $\dim(X)\in\{4,5,6,7,8\}$, we have $p_{X}=\frac{1}{3}\cdot 2+\frac{2}{3}\cdot 3=\frac{8}{3}\geq\frac{\dim(X)}{3}$, where $(p_{X})_{X\in\mathcal{L}(E)}$ is the point in $\mathcal{P}_{q}^{n}$ corresponding to $\mathcal{M}=\frac{1}{3}\mathcal{M}_{1}+\frac{2}{3}\mathcal{M}_{2}$. Hence, all subspaces of $\mathbb{F}_{q}^{n}$ are $\mu$-independent in $\mathcal{M}$, but $\mu<\frac{n}{k_{1}}$ and $n>k_{1}+k_{2}$.
Remark 6.4.
If $\mathcal{M}_{1}=\mathcal{U}_{k_{1},n}(q)$ and $\mathcal{M}_{2}=\mathcal{U}_{n,n}(q)$, then it is immediate to see that for every $\lambda\in\mathbb{Q}$ and every $k_{1}<n$, we have $\mathcal{L}(E)=\mathcal{I}_{\mu}(\mathcal{M})$, where $\mathcal{M}=(1-\lambda)\mathcal{M}_{1}+\lambda\mathcal{M}_{2}$.
6.2. Convex combination of n-2 uniform q-matroids
In this subsection, we extend previous results in the case of the convex combination of $n-2$ uniform $q$-matroids with rank values equal to $k_{i}=i+1$, for every $i\in\{1,\ldots,n-2\}$, with $n\geq 5$. Note that this cannot be done using associativity, since the convex combination of $q$-matroids is never a $q$-matroid.
For every $i\in\{1,\ldots,n-2\}$ let $\mathcal{M}_{i}=\mathcal{U}_{k_{i},n}(q)$, $\lambda_{i}=\frac{a_{i}}{b_{i}}\in\mathbb{Q}$, with for $10<\lambda_{i}<$ 1 and $\sum_{i=1}^{n-2}\lambda_{i}=1$. Let $\mathcal{M}=\sum\limits_{i=1}^{n-2}\lambda_{i}\mathcal{M}_{i}$ and let $p_{\mathcal{M}}=(p_{X})_{X\in\mathcal{L}(E)}$.
Observe that $\mathcal{I}(\mathcal{M}_{i})\subseteq\mathcal{I}(\mathcal{M}_{j})$ for every $21\leq i<j\leq n-$ 2. Moreover, it is not difficult to write down the explicit coordinates of $p_{\mathcal{M}}$. Indeed, for $X\in\mathcal{L}(E)$, we have that
$p_{X}=\begin{cases}\dim(X)&\textnormal{ if }\dim(X)\leq 2,\\ \sum\limits_{i=1}^{\dim(X)-2}(i+1)\lambda_{i}+\dim(X)\sum\limits_{i=\dim(X)-1}^{n-2}\lambda_{i}&\textnormal{ if }3\leq\dim(X)\leq n-1,\\ \sum\limits_{i=1}^{n-2}(i+1)\lambda_{i}&\textnormal{ if }\dim(X)=n.\end{cases}$
Let $\mu=\mathrm{lcm}(b_{1},\ldots,b_{n-2})$. Then $\mu$ is a denominator for $\mathcal{M}$ and clearly $\mu\geq 3$.
Theorem 6.5.
Let $\mathcal{M}=\sum\limits_{i=1}^{n-2}\lambda_{i}\mathcal{M}_{i}$, where $\mathcal{M}_{i}=\mathcal{U}_{k_{i},n}(q)$, for every $i\in\{1,\ldots,n-2\}$ let $k_{i}=i+1$ and $\lambda_{i}=\frac{a_{i}}{b_{i}}\in\mathbb{Q}$, with for $10<\lambda_{i}<$ 1 and $\sum\limits_{i=1}^{n-2}\lambda_{i}=1$. Let $\mu=\mathrm{lcm}(b_{1},\ldots,b_{n-2})$ Then all the subspaces of $\mathbb{F}_{q}^{n}$ of dimension at most $n-1$ are $\mu$-independent in $\mathcal{M}$. Moreover, if $\mu\geq\left\lceil\frac{n}{2}\right\rceil$, then $\mathcal{L}(E)=\mathcal{I}_{\mu}(\mathcal{M})$.
Proof.
It is immediate that all the subspaces in $\mathcal{L}(E)_{\leq 2}$ are $\mu$-independent, since they are strong independent. Assume that $13\leq\dim(X)\leq n-$ 1. We show that $p_{I}\geq\frac{\dim(I)}{\mu}$ for all $I\leq X$ by using induction. Assume first that $\dim(X)=3$. Then
$p_{X}=2\lambda_{1}+3(1-\lambda_{1})=3-\lambda_{1}>2,$
since $\lambda_{1}<1$. Moreover, since $\mu\geq 3$, $\frac{\dim(X)}{\mu}=\frac{3}{\mu}\leq 1$. Hence we conclude that $p_{X}\geq\frac{\dim(X)}{\mu}$. For all other $I\leq X$, this is also clear, since they are $\mu$-independent.
Assume that for all spaces $X\in\mathcal{L}(E)_{q}$, for some $13<a<n-$ 1 we have that $p_{I}\geq\frac{\dim(I)}{\mu}$ for all $I\leq X$. In particular, the following holds:
$p_{X}$ $=\sum\limits_{i=1}^{a-2}k_{i}\lambda_{i}+a\sum\limits_{i=a-1}^{n-2}\lambda_{i}$
$=\sum\limits_{i=1}^{a-2}k_{i}\lambda_{i}+a\left(1-\sum\limits_{i=1}^{a-2}\lambda_{i}\right)$
$=\sum\limits_{i=1}^{a-2}(k_{i}-a)\lambda_{i}+a\geq\frac{a}{\mu}.$
We show that also for spaces of dimension $a+1$ the same property holds. Let $Y\in\mathcal{L}(E)_{a+1}$. Then clearly for all $I\leq Y$, with $\dim(I)\leq a$, we have that $p_{I}\geq\frac{\dim(I)}{\mu}$. Moreover, we have that
$p_{Y}$ $=\sum\limits_{i=1}^{a-1}k_{i}\lambda_{i}+(a+1)\sum\limits_{i=a}^{n-2}\lambda_{i}$
$=\sum\limits_{i=1}^{a-1}k_{i}\lambda_{i}+(a+1)\left(1-\sum\limits_{i=1}^{a-1}\lambda_{i}\right)$
$=\sum\limits_{i=1}^{a-1}(k_{i}-a-1)\lambda_{i}+a+1$
$=\sum\limits_{i=1}^{a-2}(k_{i}-a-1)\lambda_{i}+(k_{a-1}-a-1)\lambda_{a-1}+a+1$
$=\sum\limits_{i=1}^{a-2}(k_{i}-a)\lambda_{i}-\sum\limits_{i=1}^{a-2}\lambda_{i}-\lambda_{a-1}+a+1$
$=\sum\limits_{i=1}^{a-2}(k_{i}-a)\lambda_{i}-\sum\limits_{i=1}^{a-1}\lambda_{i}+a+1$
$\geq\frac{a}{\mu}+1-\sum\limits_{i=1}^{a-1}\lambda_{i}$
$=\frac{a}{\mu}+\sum\limits_{i=a}^{n-2}\lambda_{i}$
$\geq\frac{a}{\mu}+\frac{1}{\mu}=\frac{a+1}{\mu}.$
This shows that every space of dimension at most $n-1$ is $\mu$-independent. We are left to show that $E$ is $\mu$-independent. Assume that $\bar{\lambda}=\min\{\lambda_{i}:i\in\{1,\dots,n-2\}\}$. Then $\bar{\lambda}=\frac{\bar{a}}{\bar{b}}$, for some $\bar{a},\bar{b}\in\mathbb{N}$. Then we have
$p_{E}$ $=\sum\limits_{i=1}^{n-2}k_{i}\lambda_{i}=\sum\limits_{i=1}^{n-2}(i+1)\lambda_{i}=\sum\limits_{i=1}^{n-2}i\lambda_{i}+1$
$\geq\sum\limits_{i=1}^{n-2}i\bar{\lambda}+1=\frac{(n-1)(n-2)}{2}\bar{\lambda}+1$
$\geq\sum\limits_{i=1}^{n-2}i\bar{\lambda}+1=\frac{(n-1)(n-2)}{2}\frac{\bar{a}}{\mu}+1$
$\geq\frac{n\bar{a}}{2\mu}+1\geq\frac{n}{2\mu}+1.$
Since $\mu\geq\left\lceil\frac{n}{2}\right\rceil$, then we conclude that $p_{E}\geq\frac{n}{\mu}$.
6.3. Cyclic flats
In this subsection we illustrate the cyclic flats of the convex combination of two uniform $q$-matroids.
Let $1<k_{1}<k_{2}<n$, $\mathcal{M}_{1}=\mathcal{U}_{k_{1},n}(q)$ and $\mathcal{M}_{2}=\mathcal{U}_{k_{2},n}(q)$. In [2, Proposition 3.30] it was shown that the cyclic flats of $\mathcal{M}_{i}$ are $\langle 0\rangle$ and $E$ and they have respectively rank values 00 and $k_{i}$, for $i=1,2$. Let $\mathcal{F}_{i}$ and $\mathcal{O}_{i}$ be respectively the collection of flats and cyclic spaces of $\mathcal{M}_{i}$, for $i=1,2$. It is clear that
$\mathcal{F}_{i}$ $=\mathcal{L}(E)_{<k_{i}}\cup\mathbb{F}_{q}^{n},$
$\mathcal{O}_{i}$ $=\mathcal{L}(E)_{>k_{i}}\cup\langle 0\rangle.$
Hence $\mathcal{F}_{1}\subseteq\mathcal{F}_{2}$ and $\mathcal{O}_{2}\subseteq\mathcal{O}_{1}$. Let $\mathcal{F}(\mathcal{M})$ and $\mathcal{O}(\mathcal{M})$ be the sets of flats and cyclic spaces of $\mathcal{M}$.
Lemma 6.6.
Let $\mathcal{M}$ be any convex combination of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. Then $\mathcal{F}(\mathcal{M})=\mathcal{F}_{2}$.
Proof.
Let $\rho_{i}$ be the rank function of $\mathcal{M}_{i}$, for $i=1,2$. From Proposition 4.4, we have that ()=$\mathcal{F}(\mathcal{M})=\{F\in\mathcal{L}(E):\operatorname{Cl}_{\rho_{1}}(F)\cap\operatorname{Cl}_{\rho_{2}}=\mathcal{L}(F)_{1}\}$. Moreover, we have $\mathcal{F}_{2}=\mathcal{F}_{1}\cup\mathcal{F}_{2}\subseteq\mathcal{F}(\mathcal{M})$. Let $X\notin\mathcal{F}_{2}$, with $\dim(X)\geq k_{2}$. Hence, for every $x\leq\mathcal{L}(E)_{1}$ we have that $\rho_{i}(X)=\rho_{i}(X+x)$ for $i=1,2$. So, we conclude that $X$ cannot lie in $\mathcal{F}(\mathcal{M})$. ∎
Lemma 6.7.
Let $\mathcal{M}$ be any convex combination of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. Then $\mathcal{O}(\mathcal{M})=\mathcal{O}_{1}$.
Proof.
Let $\rho_{i}$ be the rank function of $\mathcal{M}_{i}$, for $i=1,2$. Since $\mathcal{M}_{i}$ has no loops, from Proposition 4.6, we have that $\mathcal{O}(\mathcal{M})=\{A\in\mathcal{L}(E):\mathcal{N}_{1}(A)\cap\mathcal{N}_{2}(A)=\emptyset\}$, where $\mathcal{N}_{i}(A)=\{H\in\mathcal{H}(A):\rho_{i}(H)<\rho_{i}(A)\}$. Moreover, from Corollary 4.7, we have that $\mathcal{O}_{1}=\mathcal{O}_{1}\cup\mathcal{O}_{2}\subseteq\mathcal{O}$. Let $A\in\mathcal{L}(E)$, be a subspace with $1\leq\dim(A)\leq k_{1}$. Then $A$ is independent in both $\mathcal{M}_{i}$, since $k_{1}<k_{2}$. Then, for every $H\in\mathcal{H}(A)$, $H$ is also independent and hence $\rho_{i}(H)<\rho_{i}(A)$ and $\mathcal{N}_{1}(A)\cap\mathcal{N}_{2}(A)\neq\emptyset$. So, we conclude that $A$ cannot lie in $\mathcal{O}(A)$. ∎
Proposition 6.8.
Let $\mathcal{M}$ be any convex combination of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. The collection of cyclic flats of $\mathcal{M}$ is given by $\mathcal{Z}(\mathcal{M}):=\{\langle 0\rangle,E\}\cup\{A\in\mathcal{L}(E):k_{1}<\dim(A)<k_{2}\}$.
Proof.
The collection of cyclic flats of $\mathcal{M}$ is
$\mathcal{Z}(\mathcal{M})=\mathcal{F}(\mathcal{M})\cap\mathcal{O}(\mathcal{M})=\mathcal{F}_{2}\cap\mathcal{O}_{1}.$
7. The characteristic Puiseux polynomial
In [3, Sec. 3], the characteristic polynomial of an integer $(q,r)$-polymatroid was introduced and studied. Moreover the special case given by a $(q,1)$-polymatroid, i.e. a $q$-matroid, was then investigated in more detail in [19, Sec. 5]. In this section, we introduce a more general notion for rational $q$-polymatroids, inspired by [3]. However, this is not a polynomial anymore, but rather a truncated Puiseux series, i.e., a finite linear combination of monomials whose exponents may be rational and possibly negative. Afterwards, we discuss some properties in the context of $q$-polymatroids arising from convex combination of $q$-matroids.
Let us denote by $\mathbb{C}\{\!\{t\}\!\}$ the field of Puiseux series with coefficients in $\mathbb{C}$, that is elements of the form
$f=\sum_{k=k_{0}}^{+\infty}c_{k}t^{k/s},$
where $s$ is a positive integer and $k_{0}$ is an integer. In particular this means we allow for rational exponents of the indeterminate $t$, as long as all of these exponents have a bounded denominator and there exists a minimal one. We say that $f$ is a truncated Puiseux series if we have only finitely many terms $c_{k}\neq 0$.
The following definition is inspired by the concept of the characteristic polynomial given in [3, Def. 23].
Definition 7.1.
Let $\mathcal{M}=(\mathcal{L}(E),\rho)$ be a rational $q$-polymatroid. We define the characteristic Puiseux polynomial of $\mathcal{M}$ by
$\chi_{\mathcal{M}}(t)=\sum_{X\in\mathcal{L}(E)}\mu(\langle 0\rangle,X)t^{\ell(X)}\in\mathbb{C}\{\!\{t\}\!\},$
where $\ell(X):=\rho(E)-\rho(X)$ and $\mu$ is the Möbius function on the subspace lattice $\mathcal{L}(E)$, which is defined by
$\mu(\langle 0\rangle,X)=(-1)^{\dim(X)}q^{\dim(X)\choose 2}.$
Remark 7.2.
From Definition 7.1 we can immediately make the following observations:
  1. (a)
    The characteristic Puiseux polynomial $\chi_{\mathcal{M}}$ of any rational $q$-polymaroid $\mathcal{M}$ is indeed a truncated Puiseux series, due to the finiteness of $\mathcal{L}(E)$. Its minimal exponent is always given by zero and its maximal exponent is $\rho(E)$. Furthermore, it has only integer coefficients.
  2. (b)
    When $\mathcal{M}$ is a $(q,r)$-polymatroid for some $r\in\mathbb{N}$, the characteristic Puiseux polynomial reduces to the characteristic polynomial as defined in [3, Def. 23]. Therefore our definition is a generalization of characteristic polynomial.
  3. (c)
    One can easily verify that given two isomorphic rational $q$-polymatroids $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, their characteristic Puiseux polynomial agree, i.e., $\chi_{\mathcal{M}_{1}}(t)=\chi_{\mathcal{M}_{2}}(t)$. Note that this constitutes a generalization of the result [3, Lem. 25], since the notion of isomorphism implies the notion of lattice-equivalence, see [3, Def. 3]. Moreover the proof is identical to the one for [3, Lem. 25].
  4. (d)
    In general, every result about the characteristic polynomial of a $(q,r)$-polymatroid in [3, Sec. 3], that only uses lattice-theoretic concepts, can be done in a similar way for the characteristic Puiseux polynomial of a rational $q$-polymatroid, for instance [3, Lem. 26]. For all the results using concepts such as circuits, dependent spaces or independent spaces, we currently do not know if its possible to generalize them, using their rational $q$-polymatroid counterparts.
Now we want to examine rational $q$-polymatroids arising from rational convex combinations of two $q$-matroids. We start with an explicit example.
Example 7.3.
Let $\mathcal{M}_{1}=\mathcal{U}_{2,3}(2)$ with $q$-rank function $\rho_{1}$ and $\mathcal{M}_{2}=(\mathbb{F}_{2}^{3},\rho_{\mathcal{S}})$ be the rank-$2$ paving $q$-matroid induced by the collection
$\mathcal{S}=\bigg{\{}\bigg{\langle}\begin{matrix}0&1&0\\ 0&0&1\end{matrix}\bigg{\rangle}\bigg{\}}\subset\mathcal{L}(\mathbb{F}_{2}^{3}).$
Let $S$ be the only subspace in $\mathcal{S}$. Now we consider the rational $q$-polymatroid $\mathcal{M}=(\mathbb{F}_{2}^{3},\rho)$ arising from the convex combination of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ with rational coefficient $10<\lambda<$ 1, as in Definition 4.1. In other words $\mathcal{M}=\lambda\mathcal{M}_{1}+(1-\lambda)\mathcal{M}_{2}$. Then its characteristic Puiseux polynomial $\chi_{\mathcal{M}}$ is given by the following computation:
$\chi_{\mathcal{M}}(t)$ $=\sum_{X\in\mathcal{L}(E)}\mu(\langle 0\rangle,X)t^{\ell(X)}$
$=t^{\ell(\langle 0\rangle)}+\sum_{X\in\mathcal{L}(E)_{1}}(-1)t^{\ell(X)}+\sum_{X\in\mathcal{L}(E)_{2}\setminus\mathcal{S}}2t^{\ell(X)}+2t^{\ell(S)}-8t^{\ell(\mathbb{F}_{2}^{3})}$
$=t^{2}-7t^{1}+12t^{0}+2t^{1-\lambda}-8t^{0}$
$=t^{2}-7t+4+2t^{1-\lambda}.$
Moreover, by Remark 7.2, the characteristic polynomials $\chi_{\mathcal{M}_{1}},\chi_{\mathcal{M}_{2}}$ of the $q$-matroids $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, can also be easily computed using Definition 7.1. They are given as follows:
$\chi_{\mathcal{M}_{1}}(t)=t^{2}-7t+6\quad\text{and}\quad\chi_{\mathcal{M}_{2}}(t)=t^{2}-5t+4.$
Finally by comparing $\chi_{\mathcal{M}}$ with these characteristic polynomials we observe that
$\chi_{\mathcal{M}}(t)=\chi_{\mathcal{M}_{1}}(t)-2+2t^{1-\lambda}=\chi_{\mathcal{M}_{1}}(t)-\mu(\{0\},S)+\mu(\{0\},S)t^{1-\lambda}$
and similarly also
$\chi_{\mathcal{M}}(t)=\chi_{\mathcal{M}_{2}}(t)-2t+2t^{1-\lambda}=\chi_{\mathcal{M}_{2}}(t)-\mu(\{0\},S)t+\mu(\{0\},S)t^{1-\lambda}.$
So we see that the characteristic Puiseux polynomial of $\mathcal{M}$ can be expressed just in terms of either the characteristic polynomial of $\mathcal{M}_{1}$ or $\mathcal{M}_{2}$.
The next results generalizes the observations of Example 7.3 to arbitrary rational convex combinations of two paving $q$-matroids, arising from Proposition 5.1, having the same rank. We fix the following notation:
Let $1\leq k\leq n$ be an integer and $\mathcal{S}_{1},\mathcal{S}_{2}\subseteq\mathcal{L}(E)_{k}$ be two disjoint collections of $k$-dimensional spaces satisfying the condition from Proposition 5.1. Let $\mathcal{M}_{i}$ be the paving $q$-matroid induced by $\mathcal{S}_{i}$ and denote by $\chi_{\mathcal{M}_{i}}$ its characteristic polynomial, for $i=1,2$. Set $\mathcal{M}$ to be the convex combination of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ with rational coefficient $10<\lambda<$ 1, as in Definition 4.1.
Theorem 7.4.
The characteristic Puiseux polynomial of $\mathcal{M}$ can be expressed in terms of the characteristic polynomials of either $\mathcal{M}_{1}$ or $\mathcal{M}_{2}$. In particular we get the following formulas:
  1. (1)
    In terms of $\chi_{\mathcal{M}_{1}}$ we have:
    $\chi_{\mathcal{M}}(t)=\chi_{\mathcal{M}_{1}}(t)+(-1)^{k}q^{k\choose 2}\Big{(}|\mathcal{S}_{1}|(t^{\lambda}-t)+|\mathcal{S}_{2}|(t^{1-\lambda}-1)\Big{)}.$
  2. (2)
    In terms of $\chi_{\mathcal{M}_{2}}$ we have:
    $\chi_{\mathcal{M}}(t)=\chi_{\mathcal{M}_{2}}(t)+(-1)^{k}q^{k\choose 2}\Big{(}|\mathcal{S}_{1}|(t^{\lambda}-1)+|\mathcal{S}_{2}|(t^{1-\lambda}-t)\Big{)}.$
Proof.
Let $p_{\mathcal{M}_{1}},p_{\mathcal{M}_{2}}$ and $p_{\mathcal{M}}$ be the points in $\mathcal{P}_{q}^{n}$ corresponding to $\mathcal{M}_{1},\mathcal{M}_{2}$ and $\mathcal{M}$ respectively. Moreover, for all $X\in\mathcal{L}(E)$ we denote by $\mu(X)$ the Möbius value $\mu(\langle 0\rangle,X)$. For $i=1,2$ we have
$p_{\mathcal{M}_{i}}=(p_{\mathcal{M}_{i},X})_{X\in\mathcal{L}(E)}=\begin{cases}\;\hfil k-1\;&\text{if }X\in\mathcal{S}_{i},\\ \;\hfil\dim(X)\;&\text{if }\dim(X)\leq k,\;X\not\in\mathcal{S}_{i},\\ \;\hfil k\;&\text{otherwise}.\end{cases}$
and hence
$p_{\mathcal{M}}=(p_{\mathcal{M},X})_{X\in\mathcal{L}(E)}=\begin{cases}\;\hfil\dim(X)\;&\text{if }\dim(X)\leq k,\;X\not\in\mathcal{S}_{1}\cup\mathcal{S}_{2},\\ \;\hfil\lambda(k-1)+(1-\lambda)k\;&\text{if }X\in\mathcal{S}_{1},\\ \;\hfil\lambda k+(1-\lambda)(k-1)\;&\text{if }X\in\mathcal{S}_{2},\\ \;\hfil k\;&\text{if }\dim(X)\geq k+1.\end{cases}$
Note that $p_{\mathcal{M}_{1},E}=p_{\mathcal{M}_{2},E}=p_{\mathcal{M},E}=k$ and furthermore $p_{\mathcal{M}_{1},X}=p_{\mathcal{M}_{2},X}=p_{\mathcal{M},X}$ for all $X\in\mathcal{L}(E)\setminus(\mathcal{S}_{1}\cup\mathcal{S}_{2})$. Therefore we can split the characteristic Puiseux polynomial of $\mathcal{M}$ as follows:
$\chi_{\mathcal{M}}(t)$ $=\sum_{X\in\mathcal{L}(E)}\mu(X)t^{\ell(X)}$
$=\sum_{X\in\mathcal{L}(E)\setminus(\mathcal{S}_{1}\cup\mathcal{S}_{2})}\mu(X)t^{\ell(X)}+\sum_{X\in\mathcal{S}_{1}}\mu(X)t^{\ell(X)}+\sum_{X\in\mathcal{S}_{2}}\mu(X)t^{\ell(X)},$ (4)
where $\ell(X)=p_{\mathcal{M},E}-p_{\mathcal{M},X}$ as in Definition 7.1. Now we prove the first statement of the theorem. In Eq.(4) we observe that for all cases $X\in\mathcal{L}(E)\setminus(\mathcal{S}_{1}\cup\mathcal{S}_{2})$ we have
$\ell(X)=k-p_{\mathcal{M}_{1},X}=\ell_{1}(X),$
where $\ell_{1}$ denotes the function from Definition 7.1 corresponding to $\mathcal{M}_{1}$. For all $X\in\mathcal{S}_{1}$ we obtain
$\ell(X)=k-\lambda(k-1)-(1-\lambda)k=\lambda$
and finally for all $X\in\mathcal{S}_{2}$ we see
$\ell(X)=k-\lambda k-(1-\lambda)(k-1)=1-\lambda.$
Using these observations, we can rewrite Eq. (4) in terms of the characteristic polynomial $\chi_{\mathcal{M}_{1}}$ and further simplify this expression:
$\chi_{\mathcal{M}}(t)$ $=\chi_{\mathcal{M}_{1}}(t)-\Big{(}\sum_{X\in\mathcal{S}_{1}}\mu(X)t^{\ell_{1}(X)}+\sum_{X\in\mathcal{S}_{2}}\mu(X)t^{\ell_{1}(X)}\Big{)}+\sum_{X\in\mathcal{S}_{1}}\mu(X)t^{\lambda}+\sum_{X\in\mathcal{S}_{2}}\mu(X)t^{1-\lambda}$
$=\chi_{\mathcal{M}_{1}}(t)-(-1)^{k}q^{k\choose 2}|\mathcal{S}_{1}|t-(-1)^{k}q^{k\choose 2}|\mathcal{S}_{2}|+(-1)^{k}q^{k\choose 2}|\mathcal{S}_{1}|t^{\lambda}+(-1)^{k}q^{k\choose 2}|\mathcal{S}_{2}|t^{1-\lambda}$
$=\chi_{\mathcal{M}_{1}}(t)+(-1)^{k}q^{k\choose 2}\Big{(}|\mathcal{S}_{1}|(t^{\lambda}-t)+|\mathcal{S}_{2}|(t^{1-\lambda}-1)\Big{)},$
where we used for the second equality that $\ell_{1}(X)=1$ for all $X\in\mathcal{S}_{1}$, $\ell_{1}(X)=0$ for all $X\in\mathcal{S}_{2}$ and finally that $\mu(X)=(-1)^{k}q^{k\choose 2}$ for all $X\in\mathcal{L}(E)$. This shows the first statement of the theorem. The second statement can be deduced in a similar way by replacing $\chi_{\mathcal{M}_{1}},\ell_{1}$ with $\chi_{\mathcal{M}_{2}},\ell_{2}$ and repeat the above computation. ∎
Remark 7.5.
We want to emphasis that Theorem 7.4 does not say only that $\chi_{\mathcal{M}}$ is determined by either $\chi_{\mathcal{M}_{1}}$ or $\chi_{\mathcal{M}_{2}}$. By using [19, Thm. 5.8], we also observe that $\chi_{\mathcal{M}}$ can be computed by just using the flats of either $\mathcal{M}_{1}$ or $\mathcal{M}_{2}$. In other words, in order to determine $\chi_{\mathcal{M}}$ we do not need any concepts associated to the $q$-polymatroid $\mathcal{M}$, but rather only the well understood notion of flats of either of the $q$-matroids $\mathcal{M}_{1}$ or $\mathcal{M}_{2}$.
Note that the assumption $\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset$ is just for simplicity and Remark 7.5 still holds if we omit this condition. However, in that case the resulting characteristic Puiseux polynomial of $\mathcal{M}$ has a more intricate expression compared to Theorem 7.4.
We finally point out that it seems like situations in Theorem 7.4, in which we have a very compact expression for the characteristic Puiseux polynomial, are quite rare. Nonetheless there are still a lot of open question concerning the characteristic Puiseux polynomials of general rational $q$-polymatroids that we have not study yet. We will mention some ideas for further research in Section 9.
8. Representable q-polymatroid
In this section, we comment on the role of representable $q$-polymatroids in the polytope of all $q$-rank functions. One motivation for studying these objects is their link to rank-metric codes; see [20, 16, 24, 12].
We start by briefly recalling some basic notions on rank-metric codes; for a more detailed treatment, we refer the interested reader to [5, 8, 15]. For this purpose, we endow the space of matrices $\mathbb{F}_{q}^{n\times m}$ with the rank distance, defined by $\mathrm{d}(A,B):=\operatorname{rk}(A-B)$, for all $A,B\in\mathbb{F}_{q}^{n\times m}$.
Definition 8.1.
We say that $C\leq\mathbb{F}_{q}^{n\times m}$ is an $\mathbb{F}_{q}$-linear rank-metric code or a matrix code if $C$ is an $\mathbb{F}_{q}$-subspace of $\mathbb{F}_{q}^{n\times m}$. Its minimum distance is:
$\mathrm{d}(C):=\min\{\operatorname{rk}(M):M\in C,\;M\neq 0\}.$
We say that $C$ is an $\mathbb{F}_{q}$-$[n\times m,k,d]$ rank-metric code if it has $\mathbb{F}_{q}$-dimension $k$ and minimum distance $d$. The dual code of $C$ is defined to be $C^{\perp}=\{M\in\mathbb{F}_{q}^{n\times m}:\mathrm{Tr}(MN^{\top})=0\textnormal{ for all }N\in~{}C\}$.
The parameters $n,m,k,d$ of a rank-metric code are related by the following inequality known as Singleton bound (see e.g. [5]):
$k\leq\max\{n,m\}(\min\{n,m\}-d+1).$
Codes whose parameters meet the Singleton bound with equality are called maximum rank distance codes or MRD for short.
It is known that an $\mathbb{F}_{q}$-linear subspace of $\mathbb{F}_{q}^{n\times m}$ induces a $q$-polymatroid; see [16, 24]. One way to describe this correspondence is as follows.
Definition 8.2.
Let $C$ be an $\mathbb{F}_{q}$-$[n\times m,k]$ rank-metric code. For each subspace $X\in\mathcal{L}(E)$, we define
$C(U):=\{M\in C:\mathrm{colsp}(M)\leq U^{\perp}\}.$
It is immediate to see that for every $U\in\mathcal{L}(E)$, $C(U)$ is an $\mathbb{F}_{q}$-subspace of $\mathbb{F}_{q}^{n\times m}$. Let
$\rho:\mathcal{L}(E)\longrightarrow\mathbb{Q}_{\geq 0},\quad U\longmapsto\frac{k-\dim(C(U))}{m}.$
Then $(\mathcal{L}(E),\rho)$ is a $q$-polymatroid [16, Theorem 5.3] and we denote it by $\mathcal{M}[C]$.
A polymatroid arising from a rank-metric code is said representable. This, in particular, means that representable $q$-polymatroids correspond to rational points in the polytope $\mathcal{P}_{q}^{n}$.
Note that $\mathbb{F}_{q}^{n\times n}\cong\mathbb{F}_{q^{m}}^{n}$. If an $\mathbb{F}_{q}$-subspace of $\mathbb{F}_{q}^{n\times m}$ is in addition $\mathbb{F}_{q^{m}}$-linear, it gives rise to a $q$-matroid as follows.
Definition 8.3.
Let $C$ be a $k$-dimensional $\mathbb{F}_{q^{m}}$-linear subspace of $\mathbb{F}_{q^{m}}^{n}$. For every $W\leq\mathbb{F}_{q}^{n}$, we define
$C(W):=\{x\in C:\langle x_{1},\ldots,x_{n}\rangle_{\mathbb{F}_{q}}\leq W^{\perp}\}.$
Let $\rho:\mathcal{L}(E)\longrightarrow\mathbb{Z}_{\geq 0}$ be defined by $\rho(W):=k-\dim_{\mathbb{F}_{q^{m}}}(C(W)).$ Then $(\mathcal{L}(E),\rho)$ is a $q$-matroid [20, Thm. 24] and we also denote it by $\mathcal{M}[C]$. Note that in this case, $k$ is the dimension of $C$ over $\mathbb{F}_{q^{m}}$.
Given a rank metric code $C\leq\mathbb{F}_{q}^{n\times m}$, in general it is not immediate to determine the dimension of $C(U)$, for $U\in\mathbb{F}_{q}^{n}$. Indeed, we recall the following fact.
Proposition 8.4.
[16, Prop. 6.2] Let $C\leq\mathbb{F}_{q}^{n\times m}$ be a nonzero rank-metric code with minimum rank distance $d$ and let $d^{\perp}$ be the rank distance of $C^{\perp}$. Then for every $U\leq\mathbb{F}_{q}^{n}$,
$\rho_{\mathcal{M}[C]}(U)=\begin{cases}\frac{\dim(U)}{m}&\textnormal{ if }\dim(U)>n-d,\\ \dim(U)&\textnormal{ if }\dim(U)<d^{\perp}.\end{cases}$
There are some cases where we are able to say something more about the $q$-rank function of the $q$-polymatroid associated to a rank-metric code.
MRD codes constitute the most studied class of rank-metric codes. First of all, it is well known that MRD codes exist for every choice of parameters $n,m,d,q$. Furthermore, in [16] it is shown that if $m\geq n$, then the $q$-polymatroid arising from an MRD code is the uniform $q$-matroid of rank $n-d+1$, where $d$ is the minimum rank distance of the code. More specifically, we have the following result.
Proposition 8.5.
[16, Cor. 6.6] Let $m\geq n$ and $C\leq\mathbb{F}_{q}^{n\times m}$ be an MRD code with minimum rank distance $d$. Let $\rho_{\mathcal{M}}$ be the rank function of the $q$-polymatroid $\mathcal{M}[C]$. Then for every $A\in\mathcal{L}(E)$,
$\rho_{\mathcal{M}}(A)=\min\{n-d+1,\dim(A)\}.$
Hence $\mathcal{M}[C]$ is the uniform $q$-matroid $\mathcal{U}_{n-d+1,n}(q)$.
We can see that this is not the case for $m<n$. Indeed, we recall the following result obtained in [12].
Theorem 8.6.
[12, Theorem 3.10] Let $m<n$ and $C\leq\mathbb{F}_{q}^{n\times m}$ be an MRD code with minimum rank distance $d$. Then for every $U\leq\mathbb{F}_{q}^{n}$, we have
$\rho_{\mathcal{M}[C]}(U)=\begin{cases}\frac{n(m-d+1)}{m}&\textnormal{ if }\dim(U)\geq n-d+1,\\ \dim(U)&\textnormal{ if }\dim(U)\leq m-d+1.\end{cases}$
Moreover, $\rho_{\mathcal{M}[C]}(U)\geq\max\{1,\frac{\dim(U)}{m}\}(m-d+1)$ if $\dim(U)\in\{m-d+2,n-d\}$.
For $m=n-1$, the interval $\{m-d+2,n-d\}$ is empty for every value of the minimum distance $d$. Hence, we have the following immediate corollary.
Corollary 8.7.
Let $C\subseteq\mathbb{F}_{q}^{n\times(n-1)}$ be an MRD code with minimum rank distance $d$. Then for every $U\leq\mathbb{F}_{q}^{n}$,
$\rho_{\mathcal{M}[C]}(U)=\begin{cases}\frac{n(n-d)}{n-1}&\textnormal{ if }\dim(U)\geq n-d+1,\\ \dim(U)&\textnormal{ if }\dim(U)\leq n-d.\end{cases}$
In particular, if $d>1$, $\mathcal{M}[C]$ is not a $q$-matroid; see [12, Cor. 3.11].
In the following result, we consider two representable $q$-polymatroids arising from rank-metric codes in $\mathbb{F}_{q}^{n\times(n-1)}$. Let $C_{i}$ be an $\mathbb{F}_{q}$-$[n\times m,k_{i},d_{i}]$ MRD code, for $i=1,2$. Let $\mathcal{M}_{i}=\mathcal{M}[C_{i}]$ and let $\rho_{i}$ be its $q$-rank function, for $i=1,2$. Then, since $C_{i}$ is MRD, we have that $k_{i}=n(n-d)$ and, in particular, $\frac{k_{i}}{n}$ is an integer. The $q$-rank function $\rho_{i}$ reads as
$\rho_{i}(U)=\begin{cases}\dim(U)&\textnormal{ if }\dim(U)\leq\frac{k_{i}}{n},\\ \frac{k_{i}}{n-1}&\textnormal{ if }\dim(U)>\frac{k_{i}}{n}.\end{cases}$
The next result generalizes in some way Theorem 6.2. The proof follows the same idea as the one in Theorem 6.2, but we include it for the sake of completeness.
Theorem 8.8.
Assume that $1<k_{1}<k_{2}$ and $k_{1}+k_{2}\geq n$. Let $\lambda=\frac{a}{b}\in\mathbb{Q}$, with $a,b\in\mathbb{N}$, $\gcd(a,b)=1$ and $\mathcal{M}=(1-\lambda)\mathcal{M}_{1}+\lambda\mathcal{M}_{2}$. Then $\mu=b(n-1)$ is a denominator for $\mathcal{M}$ and $\mathcal{I}_{\mu}(\mathcal{M})=\mathcal{L}(E)$.
Proof.
Let $p_{\mathcal{M}}=(p_{X})_{X\in\mathcal{L}(E)}$ be the point in $\mathcal{P}_{q}^{n}$ corresponding to $\mathcal{M}$. Then
$p_{X}=\begin{cases}\dim(X)&\textnormal{ if }\dim(X)\leq\frac{k_{1}}{n},\\ \frac{k_{1}}{n-1}(1-\lambda)+\lambda\dim(X)&\textnormal{ if }\frac{k_{1}}{n}<\dim(X)\leq\frac{k_{2}}{n},\\ (1-\lambda)\frac{k_{1}}{n-1}+\lambda\frac{k_{2}}{n-1}&\textnormal{ if }\dim(X)>\frac{k_{2}}{n}.\end{cases}$
Let $\lambda=\frac{a}{b}\in\mathbb{Q}$, with $a,b\in\mathbb{N}$, $\gcd(a,b)=1$. Then, clearly, for $\mu=b(n-1)$ we have that $\mu p_{X}\in\mathbb{Z}$ for every $X\in\mathcal{L}(E)$. Hence, it makes sense to consider $\mu$-independent spaces. First, observe that all the subspaces $X\in\mathcal{L}(E)_{\leq\frac{k_{1}}{n}}$ are $\mu$-independent, since they are strong independent. Let $X\in\mathcal{L}(E)$ be such that $\frac{k_{1}}{n}<\dim(X)\leq\frac{k_{2}}{n}$. Then all its subspaces of dimension at most $\frac{k_{1}}{n}$ are $\mu$-independent. Let $J\leq X$ be such that $\dim(J)>\frac{k_{1}}{n}$. Then
$p_{J}$ $=\frac{k_{1}}{n-1}-\frac{ak_{1}}{b(n-1)}+\frac{a\dim(J)}{b}$
$=\frac{k_{1}b+a\dim(J)(n-1)-ak_{1}}{\mu}\geq\frac{\dim(J)}{\mu},$
where the last inequality follows from the fact that $b>a\geq 1$. Finally, we consider a subspace $J\in\mathcal{L}(E)$ with $\dim(J)>\frac{k_{2}}{n}$. Then,
$p_{J}$ $=\frac{k_{1}(b-a)}{b(n-1)}+\frac{ak_{2}}{b(n-1)}$
$=\frac{k_{1}b+a(k_{2}-k_{1})}{\mu}$
$\geq\frac{k_{2}+k_{1}(b-1)}{\mu}$
$\geq\frac{k_{1}+k_{2}}{\mu}$
$\geq\frac{n}{\mu}\geq\frac{\dim(J)}{\mu},$
where we use the fact that $b>a\geq 1$ and $n\leq k_{1}+k_{2}$. ∎
Remark 8.9.
Due to the existence of MRD codes, there are codes satisfying the assumption of Theorem 8.8, i.e. $1<k_{1}<k_{2}$ and $k_{1}+k_{2}\geq n$. For example, let $n=5$, $k_{1}=10$ and $k_{2}=15$. Then clearly $C_{1}$ is an $\mathbb{F}_{q}$-$[5\times 4,10,3]$ MRD code and $C_{2}$ is an $\mathbb{F}_{q}$-$[5\times 4,15,2]$ MRD code.
We want to conclude this section by observing that it is not possible to identify the points in $\mathcal{P}_{q}^{n}$ that correspond to representable $q$-polymatroids. This is because there are no axiomatic systems for those $q$-polymatroids. However, we might be inclined to speculate that they are vertices of $\mathcal{P}_{q}^{n}$. Unfortunately, the following example shows that this is not the case.
Example 8.10.
Let $C$ be an $\mathbb{F}_{2}$-$[3\times 2,3,2]_{2}$ MRD code and let $\mathcal{M}[C]$ be the $q$-polymatroid arising from it. Then upon fixing an ordering on $\mathcal{L}(\mathbb{F}_{q}^{n})_{i}$, for every $0\leq i\leq n$, we have that the corresponding point in $\mathcal{P}_{2}^{3}$ is $(p_{X})_{X\in\mathcal{L}(\mathbb{F}_{2}^{3})}=(0,1,1,1,1,1,1,1,3/2,3/2,3/2,3/2,3/2,3/2,3/2,3/2)$, since the $q$-rank function of $\mathcal{M}[C]$ is of the type described in Corollary 8.7. We can verify with the aid of the computer algebra system OSCAR that this is not a vertex of $\mathcal{P}_{2}^{3}$. However, there exist vertices which are representable. For example, we observe that the point
$(0,1/2,1,1,1,1/2,1,1,3/2,1,1,3/2,3/2,3/2,3/2,3/2)$
is a vertex and it can be represented as the $q$-rank function of the $q$-polymatroid arising from the $\mathbb{F}_{2}$-$[3\times 2,3,1]$ rank metric code generated by the matrices
$\begin{pmatrix}1&0\\ 0&0\\ 0&0\end{pmatrix},\quad\begin{pmatrix}0&1\\ 0&1\\ 0&0\end{pmatrix},\quad\begin{pmatrix}0&0\\ 1&1\\ 1&0\end{pmatrix}.$
9. Remarks and open questions
In this paper, we introduced and initiated the study of the polytope of all $q$-rank functions. Here, we gather some remarks and future research directions.
  1. (a)
    We highlight that the same study can be done for polymatroids. Moreover, most of our treatment could also be extended to $\mathcal{L}$polymatroids and more generally to latroids. In particular, all the arguments that depend only on the $q$-rank function and hold for a general lattice will follow in the same way.
  2. (b)
    The dimension of $\mathcal{P}_{q}^{n}$ is $\sum\limits_{i=1}^{n}\left[\begin{matrix}n\\ i\end{matrix}\right]_{q}$, making it even difficult to perform computations with the computer. Moreover, we observe that $\mathcal{P}_{q}^{n}$ contains all the $q$-rank functions of $q$-polymatroids. However, we are not able to distinguish isomorphic classes. This leads to the following question: Is it possible to find a way to identify isomorphic classes in order to reduce the dimension of $\mathcal{P}_{q}^{n}$?
  3. (c)
    Currently, we are not aware of methods for describing all the vertices of $\mathcal{P}_{q}^{n}$. In particular, while the integer vertices correspond to $q$-matroids, the rational vertices are still to be identified. We leave this investigation for future work.
  4. (d)
    By inspiration from [18] one may also study the polytope of all the nullity functions of a ($q$-)matroid. Indeed, this is non-negative, bounded, non-decreasing and supermodular.
  5. (e)
    In Section 7, we have already pointed out that the characteristic Puiseux polynomial of a rational $q$-polymatroid behaves similarly to the characteristic polynomial of a $(q,r)$-polymatroid, when it comes to the lattice-theoretic properties of this polynomial (Remark 7.2(d)). Nevertheless, the characteristic polynomial of a $(q,r)$-polymatroid has many properties in terms of independent spaces, circuits, flats, etc. So future work could be to generalize these properties to the rational $q$-polymatroid counterparts.
  6. (f)
    In Section 7 we initiated the study of the characteristic polynomial of a rational convex combination of $q$-matroids. Here, the result in Theorem 7.4 covers only a special situation. This leads to the following questions: Can we describe the characteristic Puiseux polynomial in other situations? Can we say something about its properties, in general? Is it possible to express it only in terms of the characteristic polynomials of the summands?